### 13 May 2026 v2

# Weighted algebraic topology, II

Abstract.
Extending the ‘metric spaces’ of Lawvere, we study ‘real metrics’, with values in the extended real line $\infty$$\infty$].
Formally, this ordered set is a symmetric monoidal closed category, and our structures are enriched categories on the latter. Concretely, the present goal is measuring ‘profits’ and ‘losses’ of a process, in any sense – possibly related to energy, or a variable in any science. In particular, linear real metrics derive from a potential function.
This article is Part II in a series devoted to ‘weighted algebraic topology’ – an enriched version of directed algebraic topology, where paths are measured. Part III will introduce a finer framework, more adequate to ‘quotient spaces’ (as the spheres) and better related to topology.
Key words and phrases:
Generalised metric space, directed algebraic topology, enriched categories, weighted algebraic topology
1991 Mathematics Subject Classification:
54E35, 55Uxx, 18D20
Introduction
0.1. Metrics with real values
This article is Part II of a series which began with [G3], in 2007, although at the time a sequel was not foreseen. The matter of the latter, reorganised in the general context of Directed Algebraic Topology, can also be found in Chapter 6 of the book [G4]. Both texts are freely available.
A prime structure analysed in [G3] consists of Lawvere’s generalised metrics [L1], with values in the extended positive line $\overline{\mathbb{R}}^{+}=[0,\infty]$. Here we further extend this notion, studying ‘real metrics’, with values in the extended real line $\overline{\mathbb{R}}=[-\infty,\infty]$, to express ‘profits’ and ‘losses’, in any sense – possibly related to a variable of any science, from Physics to Economy.
Two unpublished preprints of 1984 and 2015 also use metrics with values in $\overline{\mathbb{R}}$, in different perspectives. Dealing with entropy, Lawvere introduced a real valued antimetric $M(x,y)$ on the class of objects of a category $\mathbb{X}$ endowed with an ‘entropy supply’ $s\colon\mathbb{X}\to\mathbb{R}$ ([L2], p. 4). Dealing with the Legendre-Fenchel transform for real vector spaces, Willerton used real valued metrics, called $\overline{\mathbb{R}}$-metrics [Wi].
0.2. An elementary example
The present goals should be evident from the following elementary example, which does not require any prerequisite – provided we are willing to follow its concrete interpretation; definitions and results on this matter will be made precise in Section 5, after studying the general framework.
(a) We consider the vertical profile $a$ of a track, or road (in a limited area of the Earth), represented in the $xy$-plane, with points $p^{\prime}=(x^{\prime},y^{\prime})$, $p^{\prime\prime}=(x^{\prime\prime},y^{\prime\prime})$, $p_{j}=(x_{j},y_{j})$
[…]
As usual in such maps (for mountain paths, etc.), the horizontal projection on the $x$-axis represents the ‘rectified horizontal path’, that is, the rectification of the projection of the track on a horizontal plane, where the distance is measured by $\delta_{1}(p_{1},p_{2})=|x_{2}-x_{1}|$. The vertical axis measures the altitude, with distance $\delta_{2}(p,q)$ = $|y_{2}-y_{1}|$.
The distance on the path itself, and its length, are measured by the euclidean metric of the plane xyxy
(2) $\begin{array}[]{c}d(p,q)=(\delta_{1}(p,q)^{2}+\delta_{2}(p,q)^{2})^{1/2}\\[6.0pt] L_{d}(a)={\rm sup}_{\underline{p}}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;d(p_{j-1},p_{j})),\end{array}$
by a least upper bound on all finite sequences ${\underline{p}}=(p_{0},...,p_{n})$ of points of the path with increasing first coordinate.
(b) We are also interested in distinguishing ascent and descent – supposing we are going from $p^{\prime}$ to $p^{\prime\prime}$. This is not measured by the pseudometric $\delta_{2}$, but can be measured by a a real valued metric $\rho$, and its positive part $\rho^{+}=\rho{\,{\scriptstyle{\vee}}\,}0$
(3) $\rho(p_{1},p_{2})=y_{2}-y_{1},\qquad\;\;\rho^{+}(p_{1},p_{2})=\max(y_{2}-y_{1},0)\geqslant 0.$
Let us note that $\rho$ is ‘linear’ (it is finite and ‘additive’, see 2.2(c)), and expresses the increment of the gravitational potential per mass-unit, while $\rho^{+}$ is a quasi-pseudo metric [Ke] and a generalised metric in Lawvere’s sense [L1] (see 0.3). Both $\rho$ and $\rho^{+}$ are invariant up to translation and not symmetric.
(c) The real metric $\rho$ gives a real valued valuation $v_{\rho}(a)$ of the path, which, by linearity, is simply the increment in altitude from $p^{\prime}$ to $p^{\prime\prime}$
(4) $v_{\rho}(a)={\rm sup}_{\underline{p}}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\rho(p_{j-1},p_{j}))=y^{\prime\prime}-y^{\prime}\in\mathbb{R},$
with sequences $\underline{p}$} varying as above.
The total ascent $v^{+}_{\rho}(a)$ of the path is defined as its $\rho^{+}$-valuation
(5) $\begin{array}[]{c}v^{+}_{\rho}(a)=v_{\rho^{+}}(a)={\rm sup}_{p}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\rho^{+}(p_{j-1},p_{j})),\\[5.0pt] v^{+}_{\rho}(a)\,\in\,\mathbb{R}^{+}=[0,\infty[,\qquad v^{+}_{\rho}(a)\geqslant v_{\rho}(a).\end{array}$
We assume that the track is the graph of a piecewise monotone continuous function $a\colon[x^{\prime},x^{\prime\prime}]\to\mathbb{R}$. Thus $v^{+}_{\rho}(a)$ is finite: the sum of increments in altitude on the maximal intervals on which the function $a$ is increasing.
The total descent $v^{-}_{\rho}(a)$, defined as
(6) $v^{-}_{\rho}(a)=v^{+}_{\rho}(a)-v_{\rho}(a)\geqslant 0\qquad\;\;(v_{\rho}(a)=v^{+}_{\rho}(a)-v^{-}_{\rho}(a)),$
is also a real number. Owing again to the linearity of $\rho$, $v^{-}_{\rho}(a)$ can also be obtained as the total ascent of the reversed path $a^{\sharp}$, going backwards (as proved in Lemma 5.3)
(7) $v^{+}_{\rho}(a^{\sharp})={\rm sup}_{p}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\rho^{+}(p_{j},p_{j-1})).$
The absolute value $|\rho|(p_{1},p_{2})=|y_{2}-x_{2}|$ is here a pseudometric (another consequence of the linearity of $\rho$). The $|\rho|$-valuation of the path expresses the total variation of altitude
(8) $v_{|\rho|}(a)={\rm sup}_{p}(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}|\rho|(p_{j-1},p_{j})),\qquad v_{|\rho|}(a)=v^{+}_{\rho}(a)+v^{-}_{\rho}(a).$
(d) Figure (1) might also represent the value $y$ of any variable (in Physics, Chemistry, Natural Sciences, Economy, Sociology, etc.) as a function of the variable $x$ (position on a line, or time, etc.). The interpretation of $v_{\rho}(a)$, $v^{+}_{\rho}(a)$, $v^{-}_{\rho}(a)$, $v_{|\rho|}(a)$ is similar.
More complex examples, related to the gravitational or electric potential in a cartesian space, will be examined in Section 6.
0.3. Positive metrics and weighted algebraic topology
Lawvere’s article [L1] presented, in 1974, a (generalised) metric space $X$ as a small category enriched on the positive extended real line ${\bf w}^{+}=\overline{\mathbb{R}}^{+}=[0,\infty]$; the latter is viewed as a category (with morphisms $\lambda\geqslant\rho$) endowed with a strict symmetric monoidal closed structure given by the (extended) sum $\lambda+\rho$.
Concretely, $X$ is a set equipped with a (generalised) metric $\delta\colon X\!\times\!X\to{\bf w}^{+}$ (the enriched hom) satisfying the triangular inequality (enriched composition) and a second condition (enriched units)
(9) $\delta(x,y)+\delta(y,z)\geqslant\delta(x,z),\qquad\;\;0\geqslant\delta(x,x)\qquad(x,y,z\in X),$
that amounts to $\delta(x,x)=0$, for all $x\in X$. Symmetry and ‘separation’ are not required.
Similar extensions have been studied in the theory of metric spaces, like the quasi-pseudo metrics of [Ke], which take finite values, or the écarts of Bourbaki [Bk], which are symmetric.
Lawvere’s extension combines a formal elegance within category theory (one might say ‘necessity’) with concrete advantages, e.g. with respect to the existence of limits and colimits – which obviously fails for ‘finite metrics’. (An influential part of the article [L1], developed by several authors, extends part of the theory of metric spaces, like Cauchy completeness, to enriched categories.)
Breaking the symmetry property, such a metric can also express privileged directions. It was used in this sense by the present author, under the name of $\delta$-metric, as a base for an enriched version of Directed Algebraic Topology [G3, G4]. Lipschitz paths and Lipschitz homotopies are based on the standard $\delta$-interval $\delta\mathbb{I}=\delta[0,1]$, with a $\delta$-metric meant to hinder any backward move
(10) $\delta(t,t^{\prime})=t^{\prime}-t\;\mbox{ if }\>t\leqslant t^{\prime},\qquad\qquad\delta(t,t^{\prime})=\infty\;\text{ otherwise.}$
The length of a Lipschitz path $a\colon\delta\mathbb{I}\to X$ is given by the following least upper bound in $\overline{\mathbb{R}}^{+}$:
(11) $\begin{array}[]{c}L(a)={\rm sup}_{\underline{t}}\,L_{\underline{t}}(a),\\[5.0pt] L_{\underline{t}}(a)=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\delta(a(t_{j-1}),a(t_{j}))\qquad\;\;(0=t_{0}\leqslant t_{1}\leqslant...\leqslant t_{n}=1),\end{array}$
where ${\underline{t}}=(t_{j})$ stands for any finite increasing sequence in $[0,1]$.
A $\delta$-metric space $X$, or $\delta$-space, has a fundamental weighted category ${\rm w}\Pi_{1}(X)$ [G3, G4], an enrichment of the fundamental category of a directed spaces [G1, G4]: it has for objects the points of $X$, and for arrows $\xi\colon x\to x^{\prime}$ the classes of Lipschitz paths $a\colon\delta\mathbb{I}\to X$ from $x$ to $x^{\prime}$, up to the equivalence relation generated by Lipschitz homotopies with fixed endpoints. The weight of such an arrow $\xi\colon x\to x^{\prime}$ is defined as usual in a quotient, as the greatest lower bound of the length of the paths belonging to this equivalence class
(12) $w(\xi)={\rm inf}_{a\in\xi}\,L(a).$
This gives ${\rm w}\Pi_{1}(X)$ a weight, or generalised norm, in $\bf w$$^{+}$ [L1, G2, G3, G4]. All this is extended to the more general framework of ‘spaces with weighted paths’, whose finer quotients can interpret structures of Noncommutative Geometry: see [G3], Section 7, and [G4], Section 6.7.
0.4. Real valued metrics
We now replace $\bf w$$^{+}$ by a larger structure $\bf w$} based on the extended real line $\overline{\mathbb{R}}=[-\infty,\infty]$. This complete lattice is again viewed as a category, with morphisms $\lambda\geqslant\rho$, and endowed with a strict symmetric monoidal closed structure given by the extended sum $\lambda+\rho$; here we let $\infty+\infty=\infty$ (adding any element must preserve the initial object $\infty$).
A $\rho$-metric space, or $\rho$-space, is thus a small category $X$ enriched on $\bf w$}. This means a set $X$ with a mapping $\rho\colon X\!\times\!X\to\overline{\mathbb{R}}$, called a real metric, or $\rho$-metric, that satisfies the same conditions as $\delta$ in (9)
(13) $\rho(x,y)+\rho(y,z)\geqslant\rho(x,z),\qquad\;\;0\geqslant\rho(x,x)\qquad(x,y,z\in X),$
but the second condition amounts here to $\rho(x,x)=0$ or -$\infty$, for all $x\in X$ (see 2.1).
The extended real number $\rho(x,y)$ expresses a measure of the transition from $x$ to $y$. The ‘valuation’, or ‘real-valued length’, of a path is also an extended real number (see 5.2).
The basic example is now the standard $\rho$-line $\rho\mathbb{R}$, with real metric
(14) $\rho(x,y)=y-x.$
More generally, a linear $\rho$-metric on a set $X$ (see 2.2(c)) can be expressed as the variation $\rho(x,y)=\Phi(y)-\Phi(x)$ of an arbitrary function $\Phi\colon X\to\mathbb{R}$, called the potential of $\rho$ (see 3.2).
Linear real metrics are important here (and trivial in the positive case); they could also be called ‘conservative metrics’, being related to a potential. These two features will be distinguished in Part III: the more general structure of a ‘space with valued paths’ presents a linear case (possibly determined by a closed 1-form on a smooth manifold), which only corresponds to a linear $\rho$-metric in the conservative case (for an exact 1-form, or a conservative vector field on a Riemannian manifold).
0.5. Outline
The enrichments we are interested in, over $\bf w$$^{+}$ and $\bf w$}, are made clear considering the structure of a commutative quantale, in Section 1. This can be equivalently defined as a complete lattice suitably enriched, or a complete, symmetric monoidal closed category of order type.
Section 2 introduces $\rho$-metric spaces, developing their links to the positive case of $\delta$-metric spaces previously studied in [L1, G3, G4] and other works. Limits and colimits are dealt with in 2.3. The main properties, like linearity and invariance up to group actions, are outlined in 2.2 and studied in Section 3. As a critical problem, quotients of $\rho$-spaces tend to be trivial: the line $\rho\mathbb{R}$ induces a chaotic $\rho$-metric on the circle (with constant value -$\infty$), see 3.1(c).
Section 4 deals with the reflective and coreflective symmetrisations of a $\rho$-space, the associated topologies, and the associated reflective and coreflective preorders. In different situations it is convenient to investigate spaces and maps by ‘reflective’ or ‘coreflective tools’ – another critical aspect discussed below. (In Part I only the reflective procedures were needed, for $\delta$-spaces.)
Paths in a $\rho$-space and their valuation in $\overline{\mathbb{R}}$ are analysed in Section 5, extending the length of paths in a $\delta$-space. The tensor product of $\rho$-spaces is studied in Section 6, together with various examples, including real metrics associated to a conservative vector field.
0.6. Comments and sequels
(a) Directed Algebraic Topology [G4] studies ‘directed spaces’ with ‘directed algebraic structures’ produced by homotopy or homology functors: on the one hand the fundamental category (and its higher dimensional versions), on the other hand preordered homology groups. Its general aim is modelling non-reversible phenomena. It is currently used in the theory of concurrent systems, cf. [FGHMR].
Weighted Algebraic Topology, as proposed in [G3, G4], is an enrichment of this topic: the truth-valued approach of directed algebraic topology (where a path is either allowed or not) is replaced with a measure of costs, possibly infinite. The general aim is, now, measuring the cost of (possibly non-reversible) phenomena.
Here we further extend this approach, so that a cost can also be negative, expressing a gain. Various aspects of the extended theory are drastically different from the positive case, and present critical problems already mentioned in the outline, with respect to quotients and topology.
(b) The second point is the more intriguing: real metrics do have a topological content, studied throughout this paper, but a complex relation to topology. We should not expect a real metric space to have a definite topology: there is rather a range of them, from the reflective to the coreflective topology (see Section 4), and even this range can be insufficient, as in 4.8. In our perspective, the main interest of real metrics is evaluating paths (see Section 5), but different goals have already appeared in [L2, Wi].
A convincing evidence in this sense will be explored in a sequel. The Minkowski spacetime of special relativity can be given a real valued metric, derived from the metric tensor, much in the same way as a Riemannian manifold gets an ordinary metric. The path of a material or lightlike particle is continuous for the natural (euclidean!) topology, which has little to do with the metric tensor, and the associated real metric. On the other hand, these structures do evaluate the proper duration of a causal path in the spacetime.
(c) Part III will overcome these problems, and others, introducing the more flexible setting of valued spaces, or spaces with valued paths – a real valued version of the positive case already studied in [G3, G4]. In this framework an object already bears a given topology, and quotients work as one can expect: for instance, the valued line $\rm v$$\mathbb{R}$ (based on the present $\rho$-line $\rho\mathbb{R}$) has an appropriate quotient up to integral translations: the (linearly) valued circle ${\rm v}\mathbb{S}^{1}$, where a path is valued by any lifting to the line $\rho\mathbb{R}$, and a loop is valued by its (integral) winding number; this also shows that, for valued spaces, the linear case need not be conservative, as already mentioned in 0.4.
Homotopy theory for $\rho$-spaces and valued spaces will also be studied.
0.7. Notation and terminology
Open and semiopen intervals of the real line are always denoted by square brackets, like ]0,1[]0,1[, $[0,1[$, etc. The symbol $\subset$ denotes weak inclusion.
In a category, the term ‘map’ is used as a synonym of ‘morphism’; in particular, between topological spaces, it is always used in the sense of continuous mapping. In an ordered set, joins and meets are written as $x{\,{\scriptstyle{\vee}}\,}y$ and $x{\,{\scriptstyle{\wedge}}\,}y$ (also for the minimum and maximum of a pair, in a total ordering). Increasing and decreasing functions are always meant in the weak sense.
The ‘extended numbers’ $\pm\infty$ are written $\infty$ and $\overline{\infty}$}, to avoid ambiguous expressions (see 1.1(c)).
Marginal remarks and elementary proofs are written in small characters. For the sake of brevity, some technical points are referred to the preprint version of this article [G5].
0.8. Acknowledgements
I am indebted to my colleagues Ettore Carletti and Fernando Gliozzi, for long discussions on the present subject.
1. Values and weights
We begin by examining the extended real line $\overline{\mathbb{R}}=[{\overline{\infty}},\infty]$ and the extended half-line $\overline{\mathbb{R}}^{+}=[0,\infty]$, containing the values of the real or positive metrics dealt with in the next section. The terms ‘real’ and ‘positive’ will generally refer to these ordered sets.
As in Lawvere’s article [L1], based on the extended half-line, lattice terminology will refer to the natural order of $\overline{\mathbb{R}}$ and $\overline{\mathbb{R}}^{+}$, while categorical terms will refer to the direction $\lambda\geqslant\mu$ of morphisms. This (unfortunate!) reversion is required by using numbers as values of metrics; for instance, the inequality $d\geqslant d^{\prime}$ between two metrics on the same set $X$ gives a ‘comparison map’ $(X,d)\to(X,d^{\prime})$ that lifts the identity. (Dealing with antimetrics would overcome this issue, but complicate the relationship with classical metrics.)
1.1. Values, or real weights
(a) The real metrics we shall introduce take values in the extended line $\overline{\mathbb{R}}=[{\overline{\infty}},\infty]$, or – more precisely – in a structure $\bf w$} which can be analysed in two ways, based either on lattices and quantales [Mu, Ro, Ye] or on monoidal closed categories [EK, Ma, Ky]. A reader can focus on one of these aspects, but both are important: many points are more concrete in the terminology of lattices, but better motivated by category theory. (Some verifications are deferred to point (b).)
(i) As an ordered set, $\bf w$} is the extended line $\overline{\mathbb{R}}=[{\overline{\infty}},\infty]$ with the natural order $\lambda\leqslant\mu$ and the extended sum $\lambda+\mu$, where $\overline{\infty}$$\infty$$\infty$. It is a commutative unital coquantale, that is, a complete lattice with an operation $\lambda+\mu$ of abelian monoid which preserves arbitrary meets in each variable (and, in particular, the maximum $\infty$).
(ii) As a category, $\bf w$} has objects $\lambda\in\overline{\mathbb{R}}$, a unique morphism $\lambda\to\mu$ when $\lambda\geqslant\mu$ and a monoidal structure given by the extended sum $\lambda+\mu$. It is a small complete category of order type, with a symmetric monoidal closed structure.
(b) To show that $\bf w$} is indeed an ordered abelian monoid (and a symmetric monoidal category) it is sufficient to note that the additive group $\mathbb{R}$ (or any ordered abelian monoid) can be augmented adding an absorbent minimum and then an absorbent maximum.
Moreover, the fact that each mapping $\lambda+(\;)\colon{\bf w}\to{\bf w}$ preserves all meets (or colimits) is equivalent to the existence of the internal hom (computed below), by the Adjoint Functor Theorem.
(c) Joins and meets in $\bf w$} give ‘all’ limits and colimits in our category, which amount to products and sums
(15) $\begin{array}[]{llll}\text{product:}&{\,{\vee}\,}\lambda_{i},\qquad&\text{terminal object:}&{\overline{\infty}},\\[5.0pt] \text{sum:}&{\,{\wedge}\,}\lambda_{i},&\text{initial object:}&\infty.\end{array}$
The internal hom-object $\hom(\mu,\nu)\,$ is given by an extended difference, also written as $\nu-\mu$, and characterised as follows
(16) $\begin{array}[]{c}\lambda+\mu\geqslant\nu\;\iff\;\lambda\geqslant\hom(\mu,\nu),\\[5.0pt] \hom(\mu,\nu)=\nu-\mu=\min\{\lambda\in[{\overline{\infty}},\infty]\,\,|\,\lambda+\mu\geqslant\nu\}.\end{array}$
The following ‘indeterminate forms’ are thus determined:
(17) $\begin{array}[]{c}\infty-\infty=\min\{\lambda\in[{\overline{\infty}},\infty]\,\,|\,\lambda+\infty\geqslant\infty\}={\overline{\infty}},\\[5.0pt] {\overline{\infty}}-{\overline{\infty}}=\min\{\lambda\in[{\overline{\infty}},\infty]\,\,|\,\lambda+{\overline{\infty}}\geqslant{\overline{\infty}}\}={\overline{\infty}}.\end{array}$
The notation $\overline{\infty}$} allows us to distinguish the extended sum $\infty+{\overline{\infty}}=\infty$ and the extended difference $\infty-\infty={\overline{\infty}}$. Note also that $\infty$ has no additive inverse in $\bf w$}.
(d) The category $\bf w$} is *-autonomous [Ba], with dualising object the additive identity 0 and involution $\lambda^{*}=\hom(\lambda,0)=-\lambda$.
1.2. Positive values, or weights
(a) The complete lattice ${\bf w}^{+}=\overline{\mathbb{R}}^{+}=[0,\infty]$ of extended positive numbers, also called positive values, or weights, inherits a monoidal structure $\lambda+\mu$ which is also closed (in its own right). It was introduced by Lawvere [L1], in 1974 (the original notation is $\bf R$}). Weight 0 should be viewed as ‘free’, and weight $\infty$ as ‘unreachable, unaffordable, or forbidden’. Metrics with values in $\bf w$$^{+}$ will be reviewed in 2.4.
The internal hom, written here as $\hom^{+}(\mu,\nu)$, is given by the truncated difference $(\nu-\mu){\,{\scriptstyle{\vee}}\,}0$
(18) $\begin{array}[]{c}\lambda+\mu\geqslant\nu\;\iff\;\lambda\geqslant\hom^{+}(\mu,\nu),\\[5.0pt] \hom^{+}(\mu,\nu)=\min\{\lambda\in[0,\infty]\,|\,\lambda+\mu\geqslant\nu\}=(\nu-\mu){\,{\scriptstyle{\vee}}\,}0,\\[5.0pt] \hom^{+}(\infty,\infty)=\min\{\lambda\in[0,\infty]\,\,|\,\lambda+\infty\geqslant\infty\}=0.\end{array}$
Truncated difference is written $\nu-\mu$ in [L1], a notation introduced above in a different sense.
(b) The embedding ${\bf w}^{+}\subset{\bf w}$ is a strictly monoidal functor: it preserves the tensor product, meets (i.e. colimits) and non-empty joins, but not the minimum (the terminal object). It has a coreflector (right adjoint), called positive part (of a value)
(19) $(\;\;)^{+}\colon{\bf w}\to{\bf w}^{+},\qquad\;\;\lambda^{+}=\lambda{\,{\scriptstyle{\vee}}\,}0,$
which is a lax monoidal functor: it is subadditive: $(\lambda+\mu)^{+}\leqslant\lambda^{+}+\mu^{+}$, and preserves joins (i.e. limits).
(c) Families of metrics can be viewed as metrics with values in a product of the previous structures, like ${\bf w}^{p}\!\times\!({\bf w}^{+})^{q}$.
1.3. Multiplicative weights
[…]
Here we (must) take $\lambda\infty=\infty$ for all $\lambda$, including 0 (the initial object $\infty$ must be preserved).
The exponential and logarithm functions give an isomorphism (of monoidal closed categories)
(20) $\begin{array}[]{c}\exp\colon{\bf w}\;\;\raisebox{1.72218pt}{$\longrightarrow$}\hskip-15.93048pt\raisebox{-1.72218pt}{$\longleftarrow$}\;\;{\bf w}^{\scriptscriptstyle\bullet}\,\colon\!\ln,\\[5.0pt] e^{\lambda+\mu}=e^{\lambda}e^{\mu},\qquad\qquad e^{\nu-\mu}=e^{\nu}/e^{\mu}.\end{array}$
[…]
2. Real and positive metrics
Lawvere’s generalised metric spaces [L1] have been used in [G3, G4] as a first setting for weighted algebraic topology. Here we deal with the real case.
Both definitions derive from the formal theory of enriched categories [EK, Ky], which has a simple expression, by inequalities, when the base of enrichment is a category of order-type, like $\bf w$$^{+}$ or $\bf w$}. Distances in a general coquantale were introduced by R. Flagg [Fl], in 1997.
2.1. Real valued metrics
(a) A real metric space $X$, or $\rho$-metric space, or $\rho$-space, will be a set $X$ equipped with a $\rho$-metric, that is, a mapping $\rho\colon X\!\times\!X\to{\bf w}$ satisfying the axioms
(21) $\rho(x,y)+\rho(y,z)\geqslant\rho(x,z),\qquad\;\;0\geqslant\rho(x,x)\qquad(x,y,z\in X).$
We shall generally write this structure $(X,\rho)$ as $X$, and the underlying set as $|X|$; the mapping $\rho$ is written as $\rho_{X}$ when useful.
The second condition above can be rewritten as: $\rho(x,x)=0$ or $\overline{\infty}$}, since the properties $0\geqslant\rho(x,x)$ and $\rho(x,x)+\rho(x,x)\geqslant\rho(x,x)$ are only satisfied by these extended values.
The value $\overline{\infty}$} is rather exceptional in $\rho$-spaces, as shown in 3.4(c), but excluding it would make bigger problems: a chaotic $\rho$-metric only takes the value $\overline{\infty}$}, see 3.4(b).
(b) Formally, a $\rho$-space amounts to a small category $X$ enriched over the symmetric monoidal closed category $\bf w$} considered above, with a hom-value $X(x,y)$ in $\overline{\infty}$$\infty$], written as $\rho(x,y)$. This extended value will be viewed as the cost of the transition from $x$ to $y$, while $\gamma(x,y)=-\rho(x,y)$ represents the corresponding gain, or profit. (The function $\gamma$ is a ‘real antimetric’, satisfying a reverse triangular inequality.)
(c) The standard examples are related to the standard $\rho$-line $\rho\mathbb{R}$, with its cartesian and tensor powers
(22) $\rho\mathbb{R}=(\mathbb{R},\rho),\qquad\qquad\rho(x,y)=y-x.$
(d) $\rho\mathsf{Mtr}$ will denote the category of $\rho$-spaces, with cost-reducing maps $f\colon X\to Y$
(23) $\rho_{X}(x,x^{\prime})\geqslant\rho_{Y}(f(x),f(x^{\prime})),\qquad\text{for all }x,x^{\prime}\in X.$
Formally, $f$ is a w-enriched functor $X\to Y$, with hom-map $X(x,x^{\prime})\to Y(f(x),f(x^{\prime}))$.
These mappings will also be called 1-Lipschitz maps of $\rho$-metric spaces, or $\rho$-maps. Isomorphisms in this category are isometric, and will be called isometric isomorphisms or 1-Lipschitz isomorphisms. More general ‘Lipschitz maps’ will be introduced in 2.5.
The reversor endofunctor $R\colon\rho\mathsf{Mtr}\to\rho\mathsf{Mtr}$ is based on the opposite $\rho$-metric space $R(X)=X^{{\rm op}}$, with the opposite $\rho$-metric, $\rho^{{\rm op}}(x,y)=\rho(y,x)$.
2.2. An overview of the main properties
The following properties of a $\rho$-metric space $X=(X,\rho)$ will be studied below.
(a) Positivity and finiteness. A positive $\rho$-metric, or $\delta$-metric, takes values in the coquantale $\bf w$$^{+}$; this case is reviewed in 2.4. A finite $\rho$-metric takes values in $\mathbb{R}$.
(b) Flatness. We say that a point $x\in X$ is regular (resp. flat) if $\rho(x,x)=0$ (resp. $\overline{\infty}$}). The space itself is regular (resp. flat) if all its points are. Every $\rho$-space $X$ has a flat part $X_{\overline{\infty}}$, formed by its flat points. A map of $\rho$-spaces can be restricted to the flat parts.
All important examples here are regular. Flat metrics include the chaotic ones. It is easy to see that a $\rho$-space $X$ is flat if and only if $\rho(x,x^{\prime})=\pm\infty$, for all $x,x^{\prime}\in X$.
Indeed, if $X$ is flat, $\rho(x,x^{\prime})\leqslant{\overline{\infty}}+\rho(x,x^{\prime})$, whence $\rho(x,x^{\prime})=\pm\infty$. The converse is also obvious, as the only possible values of $\rho(x,x)$ are 0 and $\overline{\infty}$}.
(c) Linearity and invariance. There are two important classes of real metrics. The standard $\rho$-line $\rho\mathbb{R}$, introduced in 2.1(c), belongs to both.
First, a linear $\rho$-metric will be a finite mapping $\rho\colon X\!\times\!X\to\mathbb{R}$ that satisfies the ‘additive property’ $\rho(x,y)+\rho(y,z)=\rho(x,z)$; in particular, $\rho(x,x)=0$ for all $x$. We shall see in Section 3.2 that a linear $\rho$-metric has the form $\Delta\Phi(x,y)=\Phi(y)-\Phi(x)$, for an arbitrary ‘potential’ function $\Phi\colon X\to\mathbb{R}$. Many examples related to diverse sciences, can be obtained in this way, using for instance the gravitational potential determined by a distribution of maps or the electric potential determined by a distribution of charges, or the variation of any variable $\Phi$ (see 0.2 and 6.6).
Any chaotic metric is additive, in this sense; without the finiteness condition this is of less interest.
Secondly, a $\rho$-metric defined on an abelian group can be invariant under translations, or – equivalently – defined by a ‘real norm’ (see 3.3). Most of the metrics we shall consider on the line and its cartesian or tensor products will be of this kind.
(d) Symmetry. A symmetric $\rho$-space has $\rho=\rho^{{\rm op}}$. They form a full subcategory $\rho\mathsf{Mtr}\subset\rho\mathsf{Mtr}$, which is reflective and coreflective, see 4.1 and 4.2.
Proposition 3.6 of [G5] shows that a symmetric $\rho$-space is the sum of a positive part and a flat part (which is empty, in the important examples).
More generally, according to a general definition in Directed Algebraic Topology ([G4], 1.2.1), the $\rho$-metric space $X$ is reversive if it is isometrically isomorphic to its opposite. This is the case of $\rho\mathbb{R}$ and its cartesian or tensor products.
2.3. Comparisons, limits and colimits
(a) The notation $X\geqslant X^{\prime}$ will mean that these $\rho$-spaces have the same underlying set and $\rho_{X}\geqslant\rho_{X^{\prime}}$, or equivalently that the identity of the underlying set is a $\rho$-map $X\to X^{\prime}$, called a comparison. We also say that $\rho_{X}$ is finer than $\rho_{X^{\prime}}$.
The $\rho$-metrics on a given set form a complete lattice: given a family of them $(\rho_{i})$, their pointwise join $\rho(x,y)={\,{\vee}\,}\rho_{i}(x,y)$ is easily seen to be a $\rho$-metric.
The category $\rho\mathsf{Mtr}$ has all limits (resp. colimits), computed as in $\mathsf{Set}$ and equipped with the least (resp. greatest) $\rho$-metric that makes the structural mappings into $\rho$-maps. We write the cartesian power $(\rho\mathbb{R})^{n}$ as $\rho\mathbb{R}^{n}$.
(b) A product $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}_{i}X_{i}\,$ has the $l_{\infty}$-type $\rho$-metric
(24) $\rho({\underline{x}},{\underline{y}})={\,{\vee}\,}\rho_{i}(x_{i},y_{i})\qquad\qquad({\underline{x}}=(x_{i}),\;{\underline{y}}=(y_{i})),$
and the terminal object is the $\rho$-singleton $\rho\top$, equipped with the least $\rho$-metric, the chaotic one
(25) $\rho\top=\{*\},\qquad\rho(*,*)={\overline{\infty}}\qquad\qquad\text{(terminal object).}$
Equalisers have the restricted $\rho$-metric.
(c) A sum $\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,X_{i}$ has the obvious $\rho$-metric (and the initial object is $\varnothing$)
(26) $\rho((x,i),(y,i))=\rho_{i}(x,y),\qquad\rho((x,i),(y,j))=\infty\qquad(i\neq j).$
Coequalisers have the $\rho$-metric induced on the coequaliser in $\mathsf{Set}$, a quotient set.
Explicitly, the quotient $X/R$ of a $\rho$-metric space has the greatest $\rho$-metric making the canonical projection a $\rho$-map
(27) $\begin{array}[]{c}\rho(\xi,\eta)={\rm inf}_{\underline{x}}\;\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho(x_{2j-1},x_{2j}),\\[6.0pt] {\underline{x}}=(x_{1},...,x_{2n})\in X^{2n},\qquad\;\;x_{1}\in\xi,\;\;x_{2n}\in\eta,\\[4.0pt] x_{2j}R\,x_{2j+1}\;\text{ for }\;j=1,...,n-1.\end{array}$
2.4. Positive directed metrics
We recall the basic points of the generalised metrics introduced in [L1]. They were used as a first setting for Weighted Algebraic Topology in [G3, G4].
(a) Following the terminology that we used in [G3, G4], a $\delta$-metric space $X$, or $\delta$-space, or positive $\rho$-space, is a set $X$ equipped with a $\delta$-metric $\delta\colon X\!\times\!X\to[0,\infty]$, also called a distance, satisfying the axioms
(28) $\delta(x,y)+\delta(y,z)\geqslant\delta(x,z),\qquad\qquad 0\geqslant\delta(x,x).$
Again, $X$ is a category enriched over the symmetric monoidal closed category $\bf w$$^{+}$, with hom-value $X(x,y)=\delta(x,y)\in[0,\infty]$. We already remarked that $\delta(x,x)=0$, for every $x\in X$.
(b) $\delta\mathsf{Mtr}$ denotes the full subcategory of $\rho\mathsf{Mtr}$ of $\delta$-metric spaces. Their (cost-reducing) maps are also called 1-Lipschitz maps of $\delta$-spaces, or $\delta$-maps.
The $\delta$-metrics on a given set form a complete lattice; all limits and colimits exist in $\delta\mathsf{Mtr}$. The terminal object is now the $\delta$-singleton, equipped with the least $\delta$-metric
(29) $\delta\top=\{*\},\qquad\qquad\delta(*,*)=0.$
All non-empty products, all equalisers and all colimits are calculated as we have seen in $\rho\mathsf{Mtr}$.
(c) The embedding $\delta\mathsf{Mtr}\to\rho\mathsf{Mtr}$ has no left adjoint, as it does not preserve the terminal. The coreflector comes from the functor $(\;)^{+}\colon{\bf w}\to{\bf w}^{+}$ of (19), which truncates at 0 the negative values of a $\rho$-metric (yielding the smallest $\delta$-metric greater than the original $\rho$-metric)
(30) $\begin{array}[]{ccc}(\;\;)^{+}\colon\rho\mathsf{Mtr}\to\delta\mathsf{Mtr}&&\qquad\qquad\qquad\qquad\qquad\text{(coreflector),}\\[5.0pt] Y\mapsto Y^{+},&&\rho^{+}(y,y^{\prime})=(\rho(y,y^{\prime}))^{+}=\rho(y,y^{\prime}){\,{\scriptstyle{\vee}}\,}0,\\[5.0pt] Y^{+}\>\cong\>Y\!\times\!\delta\top,&&y\mapsto(y,*).\end{array}$
The absolute value does not give a functor from $\rho\mathsf{Mtr}$ to $\delta\mathsf{Mtr}$: for instance, it cannot work on the $\rho$-singleton, nor on any $\rho$-space having some flat point. But it does act on linear $\rho$-metrics, see 5.2(c).
(d) Again, the reversor endofunctor $R\colon\delta\mathsf{Mtr}\to\delta\mathsf{Mtr}$ is based on the opposite $\delta$-metric space $R(X)=X^{{\rm op}}$, with the opposite $\delta$-metric, $\delta^{{\rm op}}(x,y)=\delta(y,x)$.
A symmetric $\delta$-metric $\delta=\delta^{{\rm op}}$ is the same as an écart in Bourbaki [Bk]. We shall denote by $\delta\mathsf{Mtr}$ the full subcategory of $\delta\mathsf{Mtr}$ of symmetric $\delta$-metric spaces. (This category was written $\mathsf{Mtr}$ in [G3, G4], but here we want to be consistent with the parallel case $\rho\mathsf{Mtr}$.)
(e) Structures on the line and the interval. The basic example, as in [G3, G4], is the standard $\delta$-line $\delta\mathbb{R}$, with $\delta$-metric
(31) $\delta(x,y)=y-x,\;\mbox{ if }\>x\leqslant y,\qquad\qquad\delta(x,y)=\infty,\text{ otherwise,}$
where $y<x$ gives an infinite distance, with the aim of directing paths by precluding backward moves. Its cartesian powers are written $\delta\mathbb{R}^{n}$.
The sub-$\delta$-space $\delta\mathbb{I}\subset\delta\mathbb{R}$ (that is, the interval $[0,1]$ with the restricted $\delta$-metric) will play also here, as in [G3, G4]. the role of the standard interval, on which paths and homotopies are parametrised: see 5.1.
Of course we also use the euclidean line $\mathbb{R}$, with the usual euclidean metric $d$.
A different structure, used in [L1], is also of interest, the $\delta_{0}$-line $\delta_{0}\mathbb{R}$
(32) $\delta_{0}\mathbb{R}=(\rho\mathbb{R})^{+}=(\mathbb{R},\delta_{0}),\qquad\delta_{0}(x,y)=(y-x){\,{\scriptstyle{\vee}}\,}0,$
whose cartesian powers are written $\delta_{0}\mathbb{R}^{n}$.
2.5. Lipschitz maps
(a) Let $\lambda\in\mathbb{R}^{+}=[0,\infty[$. We shall write $\lambda\geqslant 0$, leaving understood that $\lambda<\infty$.
If $\lambda>0$, multiplication by $\lambda$ in $\overline{\infty}$$\infty$] preserves and reflects the order; we also assume that $0{\overline{\infty}}={\overline{\infty}}$ and $0\infty=\infty$ (cf. 1.3).
We say that a mapping $f\colon X\to Y$ between $\rho$-spaces is $\lambda$-Lipschitz, or a $\lambda$-map, if
(33) $\lambda\rho_{X}(x,x^{\prime})\geqslant\rho_{Y}(f(x),f(x^{\prime}))\qquad\qquad(x,x^{\prime}\in X),$
and we say that $f$ is a Lipschitz map, or $\rho_{\infty}$-map, if it is has some Lipschitz constant $\lambda\geqslant 0$, in the previous sense. Composing a $\lambda$-map with a consecutive $\lambda^{\prime}$-map we get a $\lambda\lambda^{\prime}$-map.
If $X$ is positive (a $\delta$-space) one can replace $\lambda$ with any $\lambda^{\prime}\geqslant\lambda$, as examined in (b). Generally, one cannot: see 2.6. (Lipschitz maps in particular cases are characterised in 3.2, 4.5, 4.6.)
The category $\rho\mathsf{Mtr}$ is a wide subcategory of the category $\rho_{\infty}\mathsf{Mtr}$ of $\rho$-spaces and Lipschitz maps. A Lipschitz isomorphism will be an isomorphism of the latter. The symmetric case is denoted as $\rho_{\infty}\mathsf{Mtr}$.
Restricting to $\delta$-metric spaces, we have the full subcategory $\delta_{\infty}\mathsf{Mtr}$ of $\delta$-metric spaces and their Lipschitz maps, also called $\delta_{\infty}$-maps, and its full subcategory $\delta_{\infty}\mathsf{Mtr}$ for the symmetric case.
(b) If $X$ is positive (a $\delta$-space) and $Y$ an arbitrary $\rho$-space, one can define, for every mapping $f\colon X\to Y$ (of sets) a Lipschitz weight, or seminorm $f||\in[0,\infty]$
(34) $||f||={\rm inf}\{\lambda\in[0,\infty[\,|\,\forall\;x,x^{\prime}\in X,\;\rho_{Y}(fx,fx^{\prime})\leqslant\lambda\delta_{X}(x,x^{\prime})\},$
[…]
The categories $\delta_{\infty}\mathsf{Mtr}$ and $\delta_{\infty}\mathsf{Mtr}$ are multiplicatively weighted (or seminormed) by the Lipschitz weight. By this we mean that every morphism $f$ is assigned a weight $||f||\in[0,\infty[$ so that
[…]
These two axioms imply that the weight of an identity can only be $1$ or 0. The wide subcategories $\delta\mathsf{Mtr}$ and $\delta\mathsf{Mtr}$ are characterised by the condition $f||\leqslant 1$ and inherit the restricted weight.
(c) The category $\rho_{\infty}\mathsf{Mtr}$ (resp. $\rho_{\infty}\mathsf{Mtr}$, $\delta_{\infty}\mathsf{Mtr}$, $\delta_{\infty}\mathsf{Mtr}$) is finitely complete and cocomplete, and the inclusion of $\rho\mathsf{Mtr}$ (resp. $\rho\mathsf{Mtr}$, $\delta\mathsf{Mtr}$, $\delta\mathsf{Mtr}$) preserves finite limits and colimits. But the ‘metric’ of a (co)limit in the larger category is only determined up to Lipschitz isomorphism.
If $X$ is a $\rho$-metric space and $\lambda\in[0,\infty[$, we shall write $\lambda X$ the same set equipped with the $\rho$-metric $\lambda\rho_{X}$. Thus, a $\lambda$-Lipschitz map is the same as a $\rho$-map $\lambda X\to Y$. This transformation preserves symmetric structures and positive structures.
2.6. Remarks
The extension to real valued metrics brings about aspects which strongly diverge from the positive case. We begin with some elementary observations. In particular, they show that the structure of $\rho\mathbb{R}$ highly restrains the Lipschitz maps defined on this space; in fact, we are more interested in maps from $\delta$-spaces to $\rho\mathbb{R}$ (as in 4.6).
(a) A $\rho_{\infty}$-map $f\colon\rho\mathbb{R}\to\rho\mathbb{R}$ is simply an affine map $f(x)=\lambda x+y_{0}$ (with $\lambda\geqslant 0$), and admits a unique Lipschitz constant, namely $\lambda$. (This will be extended to linear $\rho$-spaces, in 3.2(d).)
In fact, interchanging $x$ and $x^{\prime}$, the condition $\lambda(x^{\prime}-x)\geqslant fx^{\prime}-fx$ is equivalent to $\lambda(x^{\prime}-x)=fx^{\prime}-fx$, that is, $f(x)=\lambda x+f(0)$.
(b) Any $\lambda$-Lipschitz maps $f\colon\rho\mathbb{R}\to\delta_{0}\mathbb{R}$ is constant, and $\lambda$ is necessarily 0 (see 3.2(e)).
Letting $t>t^{\prime}$ in $\mathbb{R}$, we have $\lambda(t^{\prime}-t)\geqslant\delta_{X}(ft,ft^{\prime})\geqslant 0$, which gives $\delta_{X}(ft,ft^{\prime})=0$ for $\lambda=0$ and is impossible for $\lambda>0$.
(c) Points as maps. A point of a $\rho$-space $X$ can be identified with a $\rho$-map $\delta\top\to X$ defined on the $\delta$-singleton. Formally, the forgetful functors $\rho_{\infty}\mathsf{Mtr}\to\mathsf{Set}$ and $\rho\mathsf{Mtr}\to\mathsf{Set}$ are represented by the $\delta$-singleton (the identity of the tensor product, see 6.1). If $X$ is positive every constant mapping $f\colon X\to Y$ is a $\rho$-map, since it can be factorised as $X\to\delta\top\to Y$.
On the other hand, a $\rho_{\infty}$-map $\rho\top\to Y$ can only reach a flat point of $Y$. (The $\rho$-singleton represents the ‘flat-part functor’ mentioned in 2.2(b).)
2.7. Future and past topologies
A $\rho$-space $X=(X,\rho)$ can be equipped with several topologies, which coincide in the symmetric case.
(a) The future topology ${\rm tp}^{+}(X)$ is defined by a canonical local base at each point $x_{0}\in X$, the future open $\varepsilon$-balls
(36) $B^{+}(x_{0},\varepsilon)=\{x\in X\,|\,\rho(x_{0},x)<\varepsilon\}\qquad\;\;(\varepsilon>0\mbox{ in }\mathbb{R}),$
also written as $B^{+}_{X}(x_{0},\varepsilon)$. (Of course, this local base can be made countable.) A finer $\rho$-metric $\rho^{\prime}\geqslant\rho$ (in 2.3(a)) gives a finer future topology, with more open sets.
Dually, the past topology ${\rm tp}^{-}(X)={\rm tp}^{+}(X^{{\rm op}})$ is defined by the system of past open $\varepsilon$-balls
(37) $B-(x_{0},\varepsilon)=\{x\in X\,|\,\rho(x,x_{0})<\varepsilon\}\qquad\;\;(\varepsilon>0\mbox{ in }\mathbb{R}).$
However, we shall more often use the ‘reflective’ and ‘coreflective’ symmetric topologies, defined in 4.1 and 4.2, keeping a trace of the asymmetry of $\rho$ by a preorder (see 4.3).
(b) For a symmetric $\rho$-space $X$ the associated topological space ${\rm tp}(X)={\rm tp}^{+}(X)={\rm tp}^{-}(X)$ will generally be written $X$.
We have defined three functors
(38) $\begin{array}[]{c}{\rm tp}^{+}\colon\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top},\qquad{\rm tp}^{-}\colon\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top},\\[5.0pt] {\rm tp}\colon!\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top},\end{array}$
and will use the same notation for their restriction to the positive case.
3. Basic examples and main properties
After examining $\rho$-structures on the line and recalling $\delta$-structures on the spheres, we deal with the main properties we are interested in: linear metrics, defined by a potential function (in 3.2), and invariant metrics, defined by a real norm (in 3.3).
Several more technical properties of flat and symmetric $\rho$-spaces can be found in the preprint version of this article, Sections 3.5-3.6 [G5].
3.1. Metrics on the line and the spheres
(a) We shall generally test our notions on three basic structures, from which many others will be deduced later: the $\delta$-line $\delta\mathbb{R}$ (in (31)), the $\rho$-line $\rho\mathbb{R}$ (in 2.1(c)), and the Lawvere directed line $\delta_{0}\mathbb{R}$ (in (32)). Their future topology is determined by the following canonical local base at any point $x_{0}$:
- $\begin{array}[]{lll}\text{- for }\delta\mathbb{R}\colon&B^{+}(x_{0},\varepsilon)=[x_{0},x_{0}+\varepsilon[,\\[5.0pt] \text{- for }\rho\mathbb{R}\delta_{0}\mathbb{R}\colon&B^{+}(x_{0},\varepsilon)=\;]{\overline{\infty}},x_{0}+\varepsilon[\qquad&(\varepsilon>0).\end{array}$
This begins to make clear why we are using $\delta\mathbb{R}$ as the standard (metric) directed line: the ‘cone of the future’ at $x_{0}$ is appropriate to our goals, while that of $\delta_{0}\mathbb{R}$ is not. The $\rho$-line $\rho\mathbb{R}$ suits different goals, and its paths will be allowed to undergo losses and gains (as in the example of 0.2).
We also note that the previous metrics and topologies are invariant under translation.
(b) The orbit space $(\delta\mathbb{R})/\mathbb{Z}$ (with respect to the action of the group $\mathbb{Z}$ on the line, by integral translations) gives the standard $\delta$-circle $\delta\mathbb{S}^{1}$, where $\delta(x,y)$ is the measure of the counterclockwise arc from $x$ to $y$, with respect to the length of the circle [G3, G4]. The higher $\delta$-spheres also have an interesting structure, a quotient of the $\delta$-cube $\delta\mathbb{I}^{n}$.
(c) On the other hand, the orbit space $(\rho\mathbb{R})/\mathbb{Z}$ has the chaotic $\rho$-metric, always $\overline{\infty}$}, because two points $[x],[y]$ can be lifted to $\rho\mathbb{R}$ with arbitrarily small $\rho(x,y)$.
Relative metric spaces lack a ‘standard $\rho$-circle’. As already said in the Introduction, we shall introduce a more general setting in Part III.
As an ad hoc solution, one might introduce a notion of ‘local $\rho$-metric’, much in the same way as the chaotic preorder induced by the ordered line on the circle can be replaced by a local preorder (see [G4], 1.9.3). Yet, this complicated framework would be ineffective for other drawbacks of real metrics.
3.2. Linear real metrics and potential functions
Let $(X,\rho)$ be a $\rho$-metric space.
(a) The $\rho$-metric is linear (see 2.2) if and only if it is finite and $\rho(x,y)=-\rho(y,x)$, for all pairs $x,y$. Moreover, $\rho(x,x)=0$, for all $x$.
Indeed, if $\rho$ is linear (which includes being finite, by definition), $\rho(x,y)+\rho(y,x)=\rho(x,x)=0$. Conversely, assuming that $\rho(x,y)=-\rho(y,x)$, the triangular inequality implies the reverse one.
Let us note that the condition $\rho(x,y)=-\rho(y,x)$ does not imply that our $\rho$-metric is finite: for instance, on the set $\{0,1\}$ one can take $\rho(0,0)=0=\rho(1,1)$ and $\rho(0,1)=\infty=-\rho(1,0)$.
(b) The $\rho$-metric is linear if and only if there is a function $\Phi\colon X\to\mathbb{R}$ (a potential of $\rho$) such that $\rho$ is its variation $\Delta\Phi$
(39) $\rho(x,y)=\Delta\Phi(x,y)=\Phi(y)-\Phi(x).$
This expression plainly gives a linear $\rho$-metric. Conversely, if $\rho$ is linear, we define $\Phi(x)=\rho(x_{0},x)$, for a fixed point $x_{0}\in X$, and verify (39), using (a): $\Phi(y)-\Phi(x)=\rho(x_{0},y)-\rho(x_{0},x)=\rho(x_{0},y)+\rho(x,x_{0})=\rho(x,y).$
Concretely, $\Phi$ could be the measure of any ‘variable’ on the ‘space’ $X$. Of course, if $\Phi$ is a $C^{1}$-function on a differentiable manifold, it is the potential of its differential 1-form. One can view the variation $\rho=\Delta\Phi$ as a ‘finite differential’ of $\Phi$.
The potential $\Phi$ is determined up to an additive constant (for $X\neq\varnothing$, of course): fixing an arbitrary point $x_{0}\in X$, we can always take $\Phi(x_{0})=0$ and
(40) $\Phi(x)=\rho(x_{0},x).$
(c) Obvious examples can be deduced from the gravitational or electric fields, as in 0.2 and 6.6. A non-conservative field would produce a more general structure, studied in Part III.
A product of linear $\rho$-spaces is not linear, generally: for instance $\rho\mathbb{R}^{2}$ is not. A linear $\delta$-metric is discrete: $\delta(x,y)=0$, for all $x,y$.
(d) If $X$ and $Y$ are linear $\rho$-spaces, any $\rho$-map $f\colon X\to Y$ is an isometry. Any $\lambda$-Lipschitz map is an isometry $\lambda X\to Y$, and $\lambda$ is uniquely determined, unless $X$ has the chaotic $\delta$-metric, everywhere 0.
In fact, from $\lambda\rho_{X}(x,x^{\prime})\geqslant\rho_{Y}(fx,fx^{\prime})$ and (a) we deduce that $\lambda\rho_{X}(x,x^{\prime})$ $\rho_{Y}(fx,fx^{\prime})$, so that $f$ is an isometry $\lambda X\to Y$. Since linear $\rho$-metrics are finite, $\lambda$ is uniquely determined unless $\rho_{X}$ is everywhere 0 (which includes the case $X=\varnothing$).
(e) If $f\colon X\to Y$ is a $\lambda$-Lipschitz map from a linear $\rho$-space to a $\delta$-space, then
- $X$ has the chaotic $\delta$-metric or $\lambda=0$ (possibly both),
- $f(X)$ has the chaotic $\delta$-metric.
When these conditions are satisfied, $f$ is an arbitrary mapping.
In fact, applying (a), the condition $\lambda\rho_{X}(x,x^{\prime})\geqslant\delta_{Y}(fx,fx^{\prime})\geqslant 0$ is equivalent to $\lambda\rho_{X}(x,x^{\prime})=0=\delta_{Y}(fx,fx^{\prime})$.
3.3. Invariance and real norms
On a set $X$ acted on by a group $G$ we can consider a $\rho$-metric invariant up to the action of the group. More simply, we suppose that $X$ is an abelian group in additive notation, acting on itself.
(a) A $\rho$-metric $\rho$ on $X$ is said to be invariant up to translations if
(41) $\rho(x+z,y+z)=\rho(x,y)\qquad\qquad\qquad\qquad(x,y,z\in X).$
Here $\rho(x,x)=\rho(0,0)$, for all $x$: the metric is either regular or flat, according to the definitions of 2.2(b) – also because $X$ is non-empty. In the second (less interesting) case, we have already seen that $\rho(x,y)=\pm\infty$, for all $x,y\in X$.
(b) One can equivalently assign a real norm, or $\rho$-norm, on $X$, which means a mapping ${\underline{\mu}}\colon X\to{\bf w}$ satisfying the axioms
(42) ${\underline{\mu}}(x)+{\underline{\mu}}(y)\geqslant{\underline{\mu}}(x+y),\qquad{\underline{\mu}}(0)\leqslant 0\qquad\;\;(x,y\in X),$
where ${\underline{\mu}}(0)$ can only be 0 or $\overline{\infty}$}.
The correspondence between metric and norm is obviously expressed by the following relations:
(43) ${\underline{\mu}}(x)=\rho(0,x),\qquad\qquad\rho(x,y)={\underline{\mu}}(y-x).$
In fact, if $\rho\colon X\!\times\!X\to{\bf w}$ is an invariant $\rho$-metric, the associated mapping ${\underline{\mu}}\colon X\to{\bf w}$ satisfies (42)
$\begin{array}[]{l}{\underline{\mu}}(x)+{\underline{\mu}}(y)=\rho(0,x)+\rho(0,y)=\rho(0,x)+\rho(x,x+y)\\[3.0pt] \hskip 18.49988pt\hskip 18.49988pt\;\;\;\;\;\geqslant\;\rho(0,x+y)={\underline{\mu}}(x+y).\end{array}$
Conversely, assuming that ${\underline{\mu}}\colon X\to{\bf w}$ satisfies (42), the associated mapping $\rho\colon X\!\times\!X\to{\bf w}$ is obviously invariant up to translations and
$\rho(x,y)+\rho(y,z)={\underline{\mu}}(y-x)+{\underline{\mu}}(z-y)\,\geqslant\,{\underline{\mu}}(z-x)=\rho(x,z).$
(c) A product of abelian groups with $\rho$-metrics invariant up to translations is of the same kind. For instance all $\delta\mathbb{R}^{n}$, $\rho\mathbb{R}^{n}$, $\delta_{0}\mathbb{R}^{n}$ have a $\rho$-metric invariant up to translations.
(d) A $\rho$-metric $\rho$ on $X$ which is invariant up to translations and linear (as in $\rho\mathbb{R}$) amounts to a homomorphism of abelian groups
(44) $\Phi\colon X\to\mathbb{R},\qquad\qquad\Phi(x+y)=\Phi(x)+\Phi(y),$
which is at the same time its potential (null at $0_{X}$) and its real norm
(45) $\rho(x,y)=\Phi(y)-\Phi(x)=\Phi(y-x).$
In fact, if $\rho$ is invariant up to translations and linear, ${\underline{\mu}}(x)=\rho(0,x)=\Phi(x)$, and the relation (45) says that $\Phi$ is a homomorphism. Conversely, (45) trivially implies that $\Phi$ is a $\rho$-norm for $\rho$.
3.4. Discrete and chaotic metrics
(a) Every set $X$ can be equipped with the discrete $\rho$-metric $\rho_{D}$ (the greatest one), which is positive and symmetric
(46) $\rho_{D}(x,x)=0,\qquad\rho_{D}(x,y)=\infty\qquad\qquad(\mbox{for }x\neq y),$
giving the discrete $\rho$-metric space $\rho_{D}X$, and the left adjoint to the forgetful functor $\rho\mathsf{Mtr}\to\mathsf{Set}.$ In $\delta\mathsf{Mtr}$ these issues will be written $\delta_{D}$ and $\delta_{D}X$. (Of course the associated topology is discrete.)
A discrete $\rho$-metric space is a sum of $\delta$-singletons.
(b) Every set $X$ has a chaotic $\rho$-metric and a chaotic $\delta$-metric (the least ones), which are symmetric
(47) $\rho_{C}(x,y)={\overline{\infty}},\qquad\delta_{C}(x,y)=0\qquad\qquad(x,y\in X).$
The former gives the chaotic $\rho$-metric space $\rho_{C}X$, and the right adjoint to the forgetful functor $\rho\mathsf{Mtr}\to\mathsf{Set}$. The latter gives the chaotic $\delta$-metric space $\delta_{C}X$, and the right adjoint to $\delta\mathsf{Mtr}\to\mathsf{Set}$. (In both cases the associated topology is chaotic.)
(c) We say that a $\rho$-space is affordable if $\rho(x,x^{\prime})<\infty$, for all pairs of points. The same term applies to $\delta$-spaces.
An affordable $\rho$-space $X$ that has some points $z,z^{\prime}$ with $\rho(z,z^{\prime})={\overline{\infty}}$ is chaotic: $\rho(x,y)\leqslant\rho(x,z)+\rho(z,z^{\prime})+\rho(z^{\prime},y)={\overline{\infty}}$, for all $x,y$ (since all values are $\infty$).
4. Symmetrisation, topology and preorder
We show that the full subcategory $\rho\mathsf{Mtr}\subset\rho\mathsf{Mtr}$ of symmetric $\rho$-metric spaces (see 2.2(d)) is reflective and coreflective in $\rho\mathsf{Mtr}$, producing two associated ‘symmetric’ topologies. In the positive case, the reflector was used in [G3, G4] while the coreflector was used in [L1]. Similarly, a $\rho$-metric space has a reflective and a coreflective preorder, defined in 4.3.
Both these aspects will be of use here, in different analyses – as shown in 4.5. Their relationship to cartesian and tensor products is dealt with in Section 6.
The ‘indeterminate’ relation of real valued metrics to topology, discussed in the Introduction, is also investigated in 4.8.
4.1. The reflective symmetric metric and topology
(a) The reflector $^{\,{\scriptstyle{\wedge}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr}$ is based on the greatest symmetric $\rho$-metric $\hat{\rho}\leqslant\rho$ (already used in [G3, G4] in the positive case, and written $\delta$), which will be called here the reflective symmetric $\rho$-metric of $\rho$
(48) $\begin{array}[]{c}{}^{\,{\scriptstyle{\wedge}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr},\qquad(X,\rho)^{\,{\scriptstyle{\wedge}}\,}=(X,\hat{\rho}),\qquad\hat{\rho}=\rho{\,{\scriptstyle{\wedge}}\,}\rho^{{\rm op}},\\[5.0pt] \hat{\rho}(x,x^{\prime})={\rm inf}_{\underline{x}}(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,(\rho(x_{j-1},x_{j}){\,{\scriptstyle{\wedge}}\,}\rho(x_{j},x_{j-1}))),\\[5.0pt] \hat{\rho}(x,x^{\prime})\leqslant\rho(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\rho(x^{\prime},x),\end{array}$
where ${\underline{x}}=(x_{0},...,x_{n})$ is any finite sequence in $X$ with $x_{0}=x$ and $x_{n}=x^{\prime}$ (a step-path from $x$ to $x^{\prime}$). The unit is the comparison $X\to X^{\,{\scriptstyle{\wedge}}\,}$, the counit is the identity $Y^{\,{\scriptstyle{\wedge}}\,}=Y$, for a symmetric $\rho$-space. (The inclusion functor $\rho\mathsf{Mtr}\subset\rho\mathsf{Mtr}$ is left understood.)
Restricting to the positive case, we obtain the reflector $^{\,{\scriptstyle{\wedge}}\,}\colon\delta\mathsf{Mtr}\to\;!\delta\mathsf{Mtr}$, based on the greatest symmetric $\delta$-metric $\delta^{\,{\scriptstyle{\wedge}}\,}\leqslant\delta$. One can replace $\rho\mathsf{Mtr}$ by $\rho_{\infty}\mathsf{Mtr}$, and $\delta\mathsf{Mtr}$ by $\delta_{\infty}\mathsf{Mtr}$.
Loosely speaking, (48) ‘adjusts’ the symmetric function $\rho(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\rho(x^{\prime},x)$ to satisfy the triangular inequality in $X$. However, if this function already satisfies it, the meet $\rho{\,{\scriptstyle{\wedge}}\,}\rho^{{\rm op}}$ can be computed pointwise on $X\!\times\!X$
(49) $\hat{\rho}(x,x^{\prime})=\rho(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\rho(x^{\prime},x).$
(b) Composing with the functor ${\rm tp}\colon!\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ of (38) we have a functor $\rm tp$$^{\,{\scriptstyle{\wedge}}\,}$, also written $^{\,{\scriptstyle{\wedge}}\,}$
(50) $^{\,{\scriptstyle{\wedge}}\,}={\rm tp}^{\,{\scriptstyle{\wedge}}\,}\colon\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top}\qquad\qquad(^{\,{\scriptstyle{\wedge}}\,}\colon\rho\mathsf{Mtr}\to\mathsf{Top}).$
which equips a $\rho$-metric space with the reflective (symmetric) topology, associated to $\hat{\rho}$.
(c) For the basic examples of 3.1(a)
- $(\delta\mathbb{R})^{\,{\scriptstyle{\wedge}}\,}=\mathbb{R}$ has the euclidean metric and topology,
- $(\rho\mathbb{R})^{\,{\scriptstyle{\wedge}}\,}=\rho_{C}\mathbb{R}$ and $(\delta_{0}\mathbb{R})^{\,{\scriptstyle{\wedge}}\,}=\delta_{C}\mathbb{R}$ have a chaotic metric and chaotic topology.
In the first case, $\delta(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\delta(x^{\prime},x)=|x-x^{\prime}|$ is (already) the euclidean metric of the line. The second instance is proved in Lemma 4.7. In the third $\delta_{0}^{\,{\scriptstyle{\wedge}}\,}(x,x^{\prime})=\delta_{0}(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\delta_{0}(x^{\prime},x)=0$.
The first case above (and its consequences on cartesian powers, tensor powers and substructures) shows that the reflector is appropriate for the goals of [G3, G4]. The second shows that we cannot rely on it to get a useful topology on $\rho\mathbb{R}$.
4.2. The coreflective symmetric metric and topology
(a) The coreflector $^{\,{\scriptstyle{\vee}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr}$ is based on the least symmetric $\rho$-metric $\check{\rho}\geqslant\rho$, which will be called the coreflective symmetric $\rho$-metric of $\rho$
(51) $\begin{array}[]{ccc}{}^{\,{\scriptstyle{\vee}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr},&&(X,\rho)^{\,{\scriptstyle{\vee}}\,}=(X,\check{\rho}),\\[5.0pt] \check{\rho}=\rho{\,{\scriptstyle{\vee}}\,}\rho^{{\rm op}},&&\check{\rho}(x,x^{\prime})=\rho(x,x^{\prime}){\,{\scriptstyle{\vee}}\,}\rho(x^{\prime},x).\end{array}$
The counit is the comparison $X^{\,{\scriptstyle{\vee}}\,}\to X$.
Let us note that $\check{\rho}$ is computed pointwise on $X\!\times\!X$ because $\rho$-metrics are closed under joins. Moreover
(52) $\hat{\rho}\leqslant\rho\leqslant\check{\rho},\qquad\hat{\rho}\leqslant\rho^{{\rm op}}\leqslant\check{\rho}.$
Restricting to the positive case, we obtain the coreflector $^{\,{\scriptstyle{\vee}}\,}\colon\delta\mathsf{Mtr}\to\;!\delta\mathsf{Mtr}$, based on the least symmetric $\delta$-metric $\delta^{\,{\scriptstyle{\vee}}\,}\geqslant\delta$. Again, one can replace $\rho\mathsf{Mtr}$ by $\rho_{\infty}\mathsf{Mtr}$, and $\delta\mathsf{Mtr}$ by $\delta_{\infty}\mathsf{Mtr}$.
These right adjoints preserve products, and all limits.
(b) Composing with the functor ${\rm tp}\colon!\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ of (38) we have a functor $\rm tp$$^{\,{\scriptstyle{\vee}}\,}$, also written $^{\,{\scriptstyle{\vee}}\,}$
(53) $^{\,{\scriptstyle{\vee}}\,}={\rm tp}^{\,{\scriptstyle{\vee}}\,}\colon\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top}\qquad(^{\,{\scriptstyle{\vee}}\,}\colon\rho\mathsf{Mtr}\to\mathsf{Top}),$
which equips a $\rho$-metric space with the coreflective (symmetric) topology, associated to $\check{\rho}$. For a $\rho$-space $X$ we have the following ‘finer’ relations, for metrics and topologies:
(54) $X^{\,{\scriptstyle{\vee}}\,}\geqslant\,X,X^{{\rm op}}\,\geqslant X^{\,{\scriptstyle{\vee}}\,},\qquad\;\;X^{\,{\scriptstyle{\vee}}\,}\geqslant\,{\rm tp}^{+}X,{\rm tp}^{-}X\,\geqslant X^{\,{\scriptstyle{\vee}}\,}.$
(c) For the basic examples of 3.1(a)
- $\delta\mathbb{R}^{\,{\scriptstyle{\vee}}\,}=\delta_{D}\mathbb{R}$ has the discrete $\delta$-metric, infinite out of the diagonal, and the discrete topology,
- $\rho\mathbb{R}^{\,{\scriptstyle{\vee}}\,}$ and $\delta_{0}\mathbb{R}^{\,{\scriptstyle{\vee}}\,}$ have the euclidean metric $\check{\rho}(x,x^{\prime})=\delta_{0}^{\,{\scriptstyle{\vee}}\,}(x,x^{\prime})=|x-x^{\prime}|$ and the euclidean topology.
The coreflector is thus of interest for $\rho\mathbb{R}$, and of little use for $\delta\mathbb{R}$. We shall find convenient ways to use both the reflector and the coreflector (see 4.5); or none of them, as far as topology is concerned (in Part III).
(d) The results of this section, up to now, are abridged in the following diagrams:
[…]
Each of them contains three commutative squares, formed by full inclusions (the dashed arrows), their left adjoints (labelled with $^{\,{\scriptstyle{\wedge}}\,}$) and the right ones (with $^{\,{\scriptstyle{\vee}}\,}$).
4.3. Metrics and preorders
The asymmetry of a $\rho$-metric, overcome by symmetrisation, is also important, and can be highlighted by an associated preorder relation. Again there are two main ways: the reflective one (of interest for $\delta\mathbb{R}$, for instance) and the coreflective one (of interest for $\rho\mathbb{R}$ and $\delta_{0}\mathbb{R}$).
(a) A $\rho$-space $X$ can be equipped with the reflective preorder $x\prec_{\infty}x^{\prime}$
(56) $\begin{array}[]{c}{\rm pr}_{\infty}\colon\rho_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set},\qquad\;\;\;{\rm pr}_{\infty}(X,\rho)=(X,\prec_{\infty})\\[5.0pt] x\prec_{\infty}x^{\prime}\;\;\mbox{ if }\>\;\rho(x,x^{\prime})<\infty,\end{array}$
saying that the transition from $x$ to $x^{\prime}$ is affordable.
This functor preserves finite products. It also preserves all colimits, being left adjoint to the following embedding of preordered sets in (flat) $\rho$-spaces:
(57) $\begin{array}[]{c}{\rm mt}_{\infty}\colon{\rm p}\mathsf{Set}\to\rho_{\infty}\mathsf{Mtr},\qquad\;\;{\rm mt}_{\infty}(Y,\prec)=(Y,\rho_{\infty}),\\[5.0pt] \rho_{\infty}(y,y^{\prime})={\overline{\infty}}\;\mbox{ if }\>y\prec y^{\prime},\>\;\infty\;\text{ otherwise.}\end{array}$
For a preordered set $(Y,\prec)$, a mapping $f\colon(X,\rho)\to(Y,\rho_{\infty})$ satisfies $\rho_{\infty}(fx,fx^{\prime})\leqslant\lambda\rho(x,x^{\prime})$ (for some $\lambda\geqslant 0$) if and only if $\rho(x,x^{\prime})<\infty$ implies $\rho_{\infty}(fx,fx^{\prime})<\infty$, that is, $x\prec_{\infty}x^{\prime}$ implies $fx\prec fx^{\prime}$.
The reflective preorder is of interest for the line $\delta\mathbb{R}$, which gets the natural order $x\leqslant x^{\prime}$ (as in [G3, G4]). It is of no use for $\rho\mathbb{R}$ and $\delta_{0}\mathbb{R}$, which get the chaotic preorder.
(b) A $\rho$-space $X$ can also be equipped with the coreflective preorder $x\prec_{0}x^{\prime}$
(58) $\begin{array}[]{c}{\rm pr}_{0}\colon\rho_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set},\qquad\;\;{\rm pr}_{0}(X,\rho)=(X,\prec_{0}),\\[5.0pt] x\prec_{0}x^{\prime}\;\mbox{ if }\>\rho(x^{\prime},x)\leqslant 0,\end{array}$
saying that the backward transition, from $x^{\prime}$ to $x$, gives a gain $\gamma(x^{\prime},x)\geqslant 0$.
This functor preserves all limits; in fact it is right adjoint to a different embedding
(59) $\begin{array}[]{c}{\rm mt}_{0}\colon{\rm p}\mathsf{Set}\to\delta\mathsf{Mtr}\subset\rho_{\infty}\mathsf{Mtr},\qquad\;\;{\rm mt}_{0}(Y,\prec)=(Y,\delta_{0}),\\[5.0pt] \delta_{0}(y,y^{\prime})=0\;\mbox{ if }\>y^{\prime}\prec y,\infty\;\text{ otherwise.}\end{array}$
For a preordered set $(Y,\prec)$, a mapping $f\colon(Y,\delta_{0})\to(X,\rho)$ satisfies $\rho(fy,fy^{\prime})\leqslant\lambda\delta_{0}(y,y^{\prime})$ (for some $\lambda\geqslant 0$) if and only if $\delta_{0}(y,y^{\prime})=0$ implies $\rho(fy,fy^{\prime})=0$, that is, $y^{\prime}\prec y$ implies $fy^{\prime}\prec_{0}fy$.
For instance, the structures $\rho\mathbb{R}$ and $\delta_{0}\mathbb{R}=(\rho\mathbb{R})^{+}$ get the natural order $x\leqslant x^{\prime}$, while $\delta\mathbb{R}$ gets the discrete one.
(c) In a $\rho$-space $X$, the relation $x\prec_{0}x^{\prime}$ implies $x^{\prime}\prec_{\infty}x$ (note the reversion). In the symmetric case both preorders are equivalence relations.
4.4. Topologies and preorders on real metric lines
The previous results for the basic examples of 3.1(a) are summarised in this table
(60) $\begin{array}[]{cccccc}(X,\rho)&{\rm tp}^{+}X&\hat{\rho}(x,x^{\prime})&x\prec_{\infty}x^{\prime}&\check{\rho}(x,x^{\prime})&x\prec_{0}x^{\prime}\\[5.0pt] \delta\mathbb{R}&[x_{0},x_{0}+\varepsilon[&|x-x^{\prime}|&x\leqslant x^{\prime}&discrete&discrete\\[5.0pt] \rho\mathbb{R},\delta_{0}\mathbb{R}&[{\overline{\infty}},x_{0}+\varepsilon[&chaotic&chaotic&|x-x^{\prime}|&x\leqslant x^{\prime}\end{array}$
The future topology ${\rm tp}^{+}X$ is represented by its canonical local bases. In the last four columns we have applied the reflective and coreflective tools
(61) $\begin{array}[]{clr}{}^{\,{\scriptstyle{\wedge}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr},&\;{\rm pr}_{\infty}\colon\rho_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set}&\text{(reflective),}\\[5.0pt] {}^{\,{\scriptstyle{\vee}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr},&\;\,{\rm pr}_{0}\colon\rho_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set}\qquad&\text{(coreflective).}\end{array}$
We recall that the coreflective metric (and topology) of a $\rho$-space is always finer than the reflective one, as remarked in 4.2(b). Moreover $x\prec_{0}x^{\prime}$ implies $x^{\prime}\prec_{\infty}x$.
4.5. Reflective and coreflective analyses of maps
The Lipschitz maps between $\rho$-spaces can be studied using the previous tools.
(a) By reflective tools. The $\rho_{\infty}$-maps $f\colon\delta\mathbb{R}\to\delta\mathbb{R}$ can be characterised as the mappings $\mathbb{R}\to\mathbb{R}$ which are
(i) Lipschitz (and uniformly continuous) for the euclidean metric $d(x,x^{\prime})=|x-x^{\prime}|$,
(ii) increasing for the natural order.
The necessity of (i) and (ii) is simply proved applying to $f$ the ‘reflective tools’ of (61) (while the coreflective ones would give no information). Conversely, an increasing map $f\colon\mathbb{R}\to\mathbb{R}$ which is $\lambda$-Lipschitz for the euclidean metric is also $\lambda$-Lipschitz for $\delta\mathbb{R}$: for $x\leqslant x^{\prime}$
(62) $\lambda\delta(x,x^{\prime})=\lambda(x^{\prime}-x)=\lambda|x-x^{\prime}|\;\geqslant\;|fx-fx^{\prime}|=\delta(fx,fx^{\prime}),$
while the case $x^{\prime}<x$ is trivially satisfied, since $\lambda\delta(x,x^{\prime})=\infty$ (also for $\lambda=0$).
(b) By coreflective tools. The $\rho_{\infty}$-maps $f\colon\delta_{0}\mathbb{R}\to\delta_{0}\mathbb{R}$ are characterised in the same way, using now the coreflective metric and coreflective preorder of $\delta_{0}\mathbb{R}$, which are the usual ones on the line.
Again, we only have to check the sufficiency of conditions (i), (ii): an increasing map $f\colon\mathbb{R}\to\mathbb{R}$ which is $\lambda$-Lipschitz for the euclidean metric is also $\lambda$-Lipschitz for $\delta_{0}$: for $x\leqslant x^{\prime}$
(63) $\lambda\delta_{0}(x,x^{\prime})=\lambda(x^{\prime}-x)=\lambda|x-x^{\prime}|\geqslant|fx-fx^{\prime}|=\delta_{0}(fx,fx^{\prime}),$
while the case $x^{\prime}<x$ gives $\lambda\delta_{0}(x,x^{\prime})=0=\delta_{0}(fx,fx^{\prime})$.
(c) We have already seen that all Lipschitz maps $\rho\mathbb{R}\to\rho\mathbb{R}$ are affine (in 2.6(a)), and the Lipschitz maps $\rho\mathbb{R}\to\delta_{0}\mathbb{R}$ are constant (in 2.6(b)).
4.6. Another example
(a) A mapping $f\colon\delta\mathbb{R}\to\rho\mathbb{R}$ is $\lambda$-Lipschitz if and only if
(64) $f(x^{\prime})\leqslant f(x)+\lambda(x^{\prime}-x),\qquad\mbox{for }x\leqslant x^{\prime},$
a form of upper-right uniform continuity condition: for every $\varepsilon>0$, if $x<x_{0}+\varepsilon$ then $f(x)\leqslant f(x_{0})+\lambda\varepsilon$. Thus, the function $f$ is $\lambda$-Lipschitz for the euclidean metric on each subinterval where it is increasing, but can have any ‘downward jump’.
(b) For a mapping $f\colon\delta\mathbb{R}\to\delta_{0}\mathbb{R}$ the $\lambda$-Lipschitz condition still amounts to condition (64).
4.7. Lemma (Chaotic symmetrisations)
(a) The reflective symmetrisation of an affordable non-positive $\rho$-metric space $X$ is a chaotic $\rho$-space $\rho_{C}|X|$.
(b) In particular, $(\rho\mathbb{R})^{\,{\scriptstyle{\wedge}}\,}=\rho_{C}\mathbb{R},$ the chaotic $\rho$-line.
Proof.
In fact, assuming that $\rho(z,z^{\prime})<0$ for a pair of points, we have for all $x,x^{\prime}$ in $X$
$\hat{\rho}(x,x^{\prime})\leqslant\rho(x,z)+(2k-1)\zeta+\rho(z^{\prime},x^{\prime}),$
where $\zeta=\rho(z,z^{\prime}){\,{\scriptstyle{\wedge}}\,}\rho(z^{\prime},z)<0$ and $k$ is any integer >0>0. (We are using the step-path ${\underline{x}}=(x,z,z^{\prime},...,z,z^{\prime},x^{\prime})$, where the pair $(z,z^{\prime})$ is repeated $k$ times.) ∎
4.8. Linear metrics and topology
As remarked since the Introduction, real metrics have a complex relationship to topologies. We have so far considered four topologies on a $\rho$-metric space $X$, namely $X^{\,{\scriptstyle{\vee}}\,}\geqslant{\rm tp}^{+}X,{\rm tp}^{-}X\geqslant X^{\,{\scriptstyle{\wedge}}\,}$, but there are cases of interest where the ‘space’ already bears a natural topology, even finer than $X^{\,{\scriptstyle{\vee}}\,}$.
Extending the case of $\rho\mathbb{R}$, let us consider a topological space $X$ equipped with the linear $\rho$-metric $\rho=\Delta\Phi$ associated to a continuous potential $\Phi\colon X\to\mathbb{R}$, as in (39).
The topology of $X$ is (weakly) finer than the coreflective topology ${\rm tp}^{\,{\scriptstyle{\vee}}\,}(X,\rho)$, defined by the coreflective metric $\check{\rho}=\rho{\,{\scriptstyle{\vee}}\,}\rho^{{\rm op}}$: in fact, every open ball of the latter
(65) $B_{\Phi}(x_{0},\varepsilon)=\{x\in X\,|\,|\Phi(x)-\Phi(x_{0})|<\varepsilon\}\qquad\qquad(\varepsilon>0),$
is also open in $X$, as the $\Phi$-preimage of an open interval of $\mathbb{R}$.
These topologies coincide for $\rho\mathbb{R}$ and many related cases, but the original topology can even be strictly finer than the coreflective one: for instance, if $\Phi$ is constant, $\rho=\check{\rho}=\hat{\rho}$ is the chaotic $\delta$-metric, constant at 0, independently of the topology of $X$.
5. Paths in real valued metric spaces
As claimed in the Introduction, studying paths in a $\rho$-space $X$ is likely the best way to explore its topological information and ‘direction’. We can now make precise many topics loosely described in the elementary example of 0.2.
[…]
$X$ always denotes a $\rho$-space.
5.1. Lipschitz paths
(a) Paths (and homotopies) in $\rho$-spaces will be based on the standard $\delta$-interval $\delta\mathbb{I}$. The positivity of its metric will play an important role.
A Lipschitz path in a $\rho$-space $X$ will be a Lipschitz map $a\colon\delta\mathbb{I}\to X$. This means that there exists a (finite) $\lambda\geqslant 0$ such that
(66) $\rho_{X}(a(t),a(t^{\prime}))\leqslant\lambda(t^{\prime}-t),\qquad\mbox{for }t\leqslant t^{\prime},$
where the case $t>t^{\prime}$ is automatically satisfied. We also say that $a$ is a $\lambda$-path.
The path $a$ has a weight $||a||\in[0,\infty[$, its least Lipschitz constant: see 2.5(b).
(b) We note that
(i) the constant loop $e_{x}$ at any point $x\in X$ is Lipschitz, since $\rho(x,x)$ is 0 or $\overline{\infty}$},
(ii) Lipschitz paths are closed under concatenation: concatenating two $\lambda$-paths gives a path admitting the Lipschitz constant $2\lambda$,
(iii) Lipschitz paths are closed under (partial) reparametrisation, by any $\rho_{\infty}$-map $\varphi\colon\delta\mathbb{I}\to\delta\mathbb{I}$; in other words, if $a\colon\delta\mathbb{I}\to X$ is Lipschitz, so is obviously $a\varphi$.
(c) A Lipschitz path $\delta\mathbb{I}\to\delta\mathbb{R}$ is any increasing mapping which is Lipschitz for the euclidean metrics, as proved in 4.5(a) (for maps $\delta\mathbb{R}\to\delta\mathbb{R}$). The same applies to the reparametrisations $\varphi\colon\delta\mathbb{I}\to\delta\mathbb{I}$ considered above. We also recall that a continuous endomap of $\mathbb{I}$ need not be Lipschitz: the square root is not.
(d) For a mapping $a\colon\delta\mathbb{I}\to\rho\mathbb{R}$ the $\lambda$-Lipschitz condition amounts to
(67) $a(x^{\prime})\;\leqslant\;a(x)+\lambda(x^{\prime}-x),\qquad\;\;\mbox{for }x\leqslant x^{\prime},$
and has been described in 4.6 (for maps $\delta\mathbb{R}\to\rho\mathbb{R}$).
(e) For a mapping $a\colon\delta\mathbb{I}\to\delta_{0}\mathbb{R}$ the $\lambda$-Lipschitz condition still amounts to condition (67).
5.2. The valuation of a path
More generally, we consider now a set-theoretical path $a\colon\mathbb{I}\to X$ (a mapping of sets) in a $\rho$-space. (A restriction to Lipschitz paths $\delta\mathbb{I}\to X$ would have no relevance here.)
(a) We define the valuation $v(a)$, also written as $v_{X}(a)$ or $v_{\rho}(a)$, by the following functions:
(68) $v_{\underline{t}}(a)=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\rho(a(t_{j-1}),a(t_{j})),\qquad v(a)={\rm sup}_{\underline{t}}\,v_{\underline{t}}(a)\in{\bf w},$
where ${\underline{t}}=(t_{j})$ stands for a finite increasing sequence $0=t_{0}\leqslant t_{1}\leqslant...\leqslant t_{n}=1$. The path $a$ is said to be affordable if $v(a)<\infty$. We shall see that every Lipschitz path is affordable (Theorem 5.5(f)).
We always have $v(a)\geqslant\rho(a(0),a(1))$. If $X$ is a $\delta$-metric space, $v(a)\geqslant 0$ is also called the length of $a$ and written $L(a)$, as in [G3], Definition 1.2, and in [G4], 6.1.8.
(b) We define the upper valuation, or total ascent, of $a$ by its length in the $\delta$-metric space $X^{+}$
(69) $\begin{array}[]{c}v^{+}(a)=L_{X^{+}}(a)={\rm sup}_{\underline{t}}\,\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\rho^{+}(a(t_{j-1}),a(t_{j}))\in{\bf w}^{+},\\[5.0pt] v^{+}(a)\,\geqslant\,v(a).\end{array}$
If $v(a)<\infty$, we also define the lower valuation, or total descent, of $a$ as the extended difference
(70) $v^{-}(a)=v^{+}(a)-v(a)\in{\bf w}^{+},$
which is $\infty$ when $v^{+}(a)=\infty$ or $v(a)={\overline{\infty}}$ (possibly both). Therefore
(71) $v(a)=v^{+}(a)-v^{-}(a)\qquad\qquad(\text{if }v(a){\,{\scriptstyle{\vee}}\,}v^{+}(a)<\infty).$
If $X$ is a $\delta$-space, we are only interested in its length $L(a)=v(a)=v^{+}(a)\in{\bf w}^{+}$. (Here $v^{-}(a)=0$, provided that $v(a)<\infty$.)
(c) Absolute value. If $\rho$ is linear, its absolute value $|\rho|(x,y)=|\rho(x,y)|$ is easily seen to be a finite symmetric $\delta$-metric (generally called a pseudometric in the literature).
The total variation of the path $a$ will be its $|\rho|$-valuation
(72) $\begin{array}[]{l}L_{|\rho|}(a)={\rm sup}_{\underline{x}}\,\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,|\rho|(x_{j-1},x_{j})={\rm sup}_{\underline{x}}\,\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,|(\Phi(x_{j})-\Phi(x_{j-1})|\\[5.0pt] \qquad\;\;\;\;=v^{+}_{\rho}(a)+v^{-}_{\rho}(a).\end{array}$
5.3. Lemma
If $\rho$ is linear, with potential $\Phi$, for any path $a\colon\mathbb{I}\to X$ (with reversed path $a^{\sharp}$)
(73) $\begin{array}[]{c}v(a)=\rho(a(0),a(1))=\Phi(a(1))-\Phi(a(0))\in\mathbb{R},\\[3.0pt] v(a^{\sharp})=-v(a),\qquad\qquad v^{-}(a)=v^{+}(a^{\sharp}).\end{array}$
Proof.
The first relation is obvious and the second is a consequence. As to to the third, we recall that $\rho(y,x)=-\rho(x,y)$ and note that, for $\lambda\in\mathbb{R}$, we have $\lambda=\lambda^{+}-(-\lambda)^{+}$.
Letting $\underline{t}$} be any finite increasing sequence of points in $\mathbb{I}$, with $0=t_{0}\leqslant t_{1}\leqslant...\leqslant t_{n}=1$ and $p_{j}=a(t_{j})$, we have by linearity the following relations:
$\begin{array}[]{c}v(a)=\rho(p_{0},p_{1})=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho(p_{j-1},p_{j})=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho^{+}(p_{j-1},p_{j})\!-\!\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho^{+}(p_{j},p_{j-1}),\\[5.0pt] \raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho^{+}(p_{j-1},p_{j})=v(a)+\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho^{+}(p_{j},p_{j-1}),\\[5.0pt] v^{+}(a)=v(a)+{\rm sup}_{\underline{t}}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho^{+}(p_{j},p_{j-1})))=v(a)+v^{+}(a^{\sharp}),\end{array}$
so that $v^{+}(a^{\sharp})$ is indeed $v^{+}(a)-v(a)=v^{-}(a)$. ∎
5.4. Examples and remarks
(a) For a path $a$ in $\delta\mathbb{R}$, L(a)=a(1)a(0)L(a)=a(1)-a(0) if $a$ is increasing, and $L(a)=\infty$ otherwise.
(b) For a path $a$ in the linear $\rho$-space $\rho\mathbb{R}$
- v(a)=a(1)a(0)v(a)=a(1)-a(0) is always finite, but $v^{+}(a)$ and $v^{-}(a)$ can also be infinite (for instance, one can take the projection on a line of a Peano curve in the square),
- also the weight $a||$ (of $a\colon\delta\mathbb{I}\to\rho\mathbb{R}$) can be infinite: a classical example is $a(t)=\sqrt{t}$,
- if $a$ is increasing, $v^{+}(a)=v(a)=a(1)-a(0)$ and $v^{-}(a)=0$,
- if $a$ is decreasing, $v^{-}(a)=-v(a)=a(0)-a(1)$ and $v^{+}(a)=0$.
(c) For a path $a$ in $\delta_{0}\mathbb{R}$, the length $L(a)$ is the upper valuation of $a$ in $\rho\mathbb{R}$ and can be infinite. For an increasing (resp. decreasing) path, L(a)=a(1)a(0)L(a)=a(1)-a(0) (resp. $L(a)=0$).
(d) For a (set-theoretical) path $a$ in a discrete $\rho$-space $\rho_{D}X=\delta_{D}X$, $L(a)$ is 0 or $\infty$, according to the fact that $a$ is constant or not.
In a chaotic $\rho$-space $\rho_{C}X$, $v(a)={\overline{\infty}}$. In a chaotic $\delta$-space $\delta_{C}X$, $L(a)=0$.
5.5. Theorem (Path valuation)
Let $X$ be a $\rho$-metric space.
The valuation of paths satisfies the following properties, where $a\colon\mathbb{I}\to X$ is a set-theoretical path, $e_{x}$ is the constant path at $x\in X$, aba*b is the concatenation of two consecutive paths, $\varphi\colon\mathbb{I}\to\mathbb{I}$ is a partial reparametrisation (an increasing map) and $a||\leqslant\infty$ is the Lipschitz weight of $a\colon\mathbb{I}\to X$ as a mapping defined on the $\delta$-space $\delta\mathbb{I}$.
[…]
Proof.
There is a similar result in [G4], 6.1.8, for the positive case. Note that we can always replace a partition $\underline{t}$} by a finer one $\underline{t}$$^{\prime}$, as the triangular property gives $v_{\underline{t}}(a)\leqslant v_{{\underline{t}}^{\prime}}(a)$.
Point (a) is obvious. For (b), the inequality $v(a*b)\leqslant v(a)+v(b)$ follows easily from the previous remark: given a partition $\underline{t}$} for c=abc=a*b, call $\underline{t}$$^{\prime}$ its refinement by introducing the point $t=1/2$ (if missing); then $v_{\underline{t}}(c)\leqslant v_{{\underline{t}}^{\prime}}(c)\leqslant v(a)+v(b)$. For the other inequality, it is sufficient to note that a partition for $a$ and one for $b$ yield a partition of $[0,2]$, which can be scaled down to the standard interval.
Point (e) follows from the fact that we are computing the mapping $v$ on the same sequences: if $(t_{j})$ is an increasing sequence in $\mathbb{I}$, $(\varphi(t_{j}))$ is also; conversely, given $(t_{j})$, there is an increasing sequence $(s_{j})$ such that $\varphi(s_{j})=t_{j}$, for all indices $j$. (Note also that $\varphi(0)=0$ and $\varphi(1)=1$.)
[…]
Finally, points (f) and (g) are obvious, because
[…]
5.6. Spaces with distinguished paths
This point is addressed to a reader with a basic knowledge of ‘spaces with directed paths’, or d-spaces, our main structure for Directed Algebraic Topology, introduced in [G1], extensively studied in [G4] and frequently used in the theory of concurrency (see [FGHMR] and its references).
In [G3, G4] a $\delta$-space $X$ has an associated d-space ${\bf d}X$ on the same set, equipped with the reflective topology defined in 4.1, and directed paths consisting of all continuous mappings $a\colon\mathbb{I}\to X$ which are affordable, in the sense that $v(a)<\infty$; in other words, their length $L(a)$ is finite.
This can be extended to $\rho$-spaces (still using the reflective topology): the axioms of $\delta$-spaces are still satisfied (by Theorem 5.5(a), (b), (d)), and $\rho$-maps preserve the directed paths (by 5.5(g)).
Yet, we have already seen that this extension cannot be satisfactory: $\rho\mathbb{R}$ and several related $\rho$-spaces would get the chaotic topology. Using the final topology on $X$ for the affordable paths might be preferable, but is not an extension of the positive case, and would give strange topologies, of little interest, to basic spaces like all $\delta\mathbb{R}^{n}$, for $n>1$.
Essentially, this problem arises because we are working with ‘directed metrics’: if $X$ is a first-countable, locally path connected space, the final topology for all its (continuous) maps $\mathbb{I}\to X$ is precisely the original topology, as proved in [ChSW], Proposition 3.11, in an article about diffeological spaces.
Finally, this is another evidence of the vague relationship of real metrics to topology. As claimed in the Introduction, Part III will give a better solution to all these problems.
6. Tensor products, other examples
The category $\rho\mathsf{Mtr}$ has a ‘natural’ symmetric monoidal closed structure, that extends the structure of $\delta\mathsf{Mtr}$ [L1, G3, G4], with the same formulas. We also examine the relationship between the tensor product and the cartesian product. The elementary example of the Introduction is reviewed in 6.5; the gravitational and elastic potential in 6.6, 6.7.
6.1. Tensor product
(a) The tensor product $X\otimes Y$ of $\rho$-spaces is the cartesian product of the underlying sets, endowed with the $l_{1}$-type $\rho$-metric (instead of the $l_{\infty}$-type $\rho$-metric of the categorical product)
(74) $\rho((x,y),(x^{\prime},y^{\prime}))=\rho_{X}(x,x^{\prime})+\rho_{Y}(y,y^{\prime}).$
The unit is the $\delta$-singleton $\delta\top$ (see 2.4(b)). On the other hand, $X\otimes\rho\top=\rho_{C}|X|$, for all affordable $\rho$-spaces $X$.
This construction solves the usual universal problem, with respect to mappings which are 1-Lipschitz in each variable. The exponential $Z^{Y}$ is the set of 1-Lipschitz maps $Y\to Z$ equipped with the $\rho$-metric
(75) $\rho(h,k)={\rm sup}_{y\in Y}\,\rho_{Z}(h(y),k(y))\qquad\;\;(h,k\colon Y\to Z\mbox{ in }\rho\mathsf{Mtr}).$
The proof of the adjunction is standard, and can be found in point (d). The tensor powers of the basic $\rho$-structures of the line are written as $\delta\mathbb{R}^{\otimes n}$, $\rho\mathbb{R}^{\otimes n}$ and $\delta_{0}\mathbb{R}^{\otimes n}$.
(b) Restricting to positive $\delta$-metrics, the cartesian and tensor product satisfy the following inequalities:
(76) $X\!\times\!Y\leqslant X\otimes Y\leqslant 2(X\!\times\!Y)\qquad\qquad(\text{in }\delta\mathsf{Mtr}).$
In $\delta_{\infty}\mathsf{Mtr}$ these products are thus isomorphic, and denote isomorphic functors in two variables. This is not true in $\delta\mathsf{Mtr}$, and we shall always distinguish such objects (and functors): the notation $X\!\times\!Y$ (resp. $X\otimes Y$) will always denote the structure given by the $l_{\infty}$-type (resp. $l_{1}$-type) $\delta$-metric.
(c) For $\rho$-metrics we still have
(77) $X\otimes Y\leqslant 2(X\!\times\!Y)\qquad\qquad(\text{in }\rho\mathsf{Mtr}),$
but the first inequality in (76) fails (also for symmetric $\rho$-spaces). For instance, the $\rho$-spaces $\delta\top\!\times\!\rho\top=\delta\top$ and $\delta\top\otimes\rho\top=\rho\top$ give the reverse inequality, and are not even Lipschitz-isomorphic.
(d) Finally, to prove the adjunction, we verify the bijective correspondence $f\leftrightarrow g$ between 1-Lipschitz maps
(78) $f\colon X\otimes Y\to Z,\qquad g\colon X\to Z^{Y}\qquad\qquad(\text{in }\rho\mathsf{Mtr}),$
linked by the relation $f(x,y)=g(x)(y)$, for all $x\in X$, $y\in Y$. (These quantifiers are left understood, below).
Given $f$, the relation $\rho_{Z}(f(x,y),f(x^{\prime},y^{\prime}))\leqslant\rho_{X}(x,x^{\prime})+\rho_{Y}(y,y^{\prime})$ shows that
- every map $g(x)=f(x,-)\colon Y\to Z$ is 1-Lipschitz, because (taking into account that $\rho_{X}(x,x)\leqslant 0$)
$\rho_{Z}(f(x,y),f(x,y^{\prime}))\leqslant\rho_{X}(x,x)+\rho_{Y}(y,y^{\prime})\leqslant\rho_{Y}(y,y^{\prime}),$
- the map $g\colon X\to Z^{Y}$ is 1-Lipschitz, because we have, in the same way
$\begin{array}[]{c}\rho_{Z}(g(x)(y),g(x^{\prime})(y))=\rho_{Z}(f(x,y),f(x^{\prime},y))\\[3.0pt] \qquad\qquad\qquad\qquad\leqslant\;\rho_{X}(x,x^{\prime})+\rho_{Y}(y,y)\leqslant\rho_{X}(x,x^{\prime}).\end{array}$
The other implication is similarly proved, inserting a point $(x^{\prime},y)$ between $(x,y)$ and $(x^{\prime},y^{\prime})$.
6.2. Theorem (Properties of the tensor product)
Let $X,Y$ be $\rho$-metric spaces.
(a) The reflector $^{\,{\scriptstyle{\wedge}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr}$ is a strict monoidal functor
(79) $(X\otimes Y)^{\,{\scriptstyle{\wedge}}\,}=X^{\,{\scriptstyle{\wedge}}\,}\otimes Y^{\,{\scriptstyle{\wedge}}\,}$,
and forms, with the embedding $\rho\mathsf{Mtr}\subset\rho\mathsf{Mtr}$, a strict adjunction of monoidal categories.
(b) The coreflector $^{\,{\scriptstyle{\vee}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr}$ is a lax monoidal functor
(80) $(X\otimes Y)^{\,{\scriptstyle{\vee}}\,}\leqslant X^{\,{\scriptstyle{\vee}}\,}\otimes Y^{\,{\scriptstyle{\vee}}\,}$,
with comparison $X^{\,{\scriptstyle{\vee}}\,}\otimes Y^{\,{\scriptstyle{\vee}}\,}\to(X\otimes Y)^{\,{\scriptstyle{\vee}}\,}$. It forms a strict/lax adjunction.
(c) As to the reflective and coreflective preorders (in 4.3)
[…]
(82) $(x\prec_{0}x^{\prime}\mbox{ in }X,\;y\prec_{0}y^{\prime}\mbox{ in }Y)\implies(x,y)\prec_{0}(x^{\prime},y^{\prime})\mbox{ in }X\otimes Y.$
(d) For a set-theoretical path $\langle a,b\rangle\colon\mathbb{I}\to X\otimes Y$
(83) $v(\langle a,b\rangle)=v(a)+v(b).$
Proof.
We replace the pair $X,Y$ with a finite family of $\rho$-metric spaces $X_{1},...,X_{n}$; this will give a simpler notation and a clearer proof. An element of the set $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}_{i}|X_{i}|$ is written as ${\underline{x}}=(x_{i})$. We use the notation $\rho_{X}(x,x^{\prime})=X(x,x^{\prime})$, which comes from viewing a $\rho$-metric space as an enriched category (in 2.1).
(a) We now prove that $(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}_{i}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}=\;\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}_{i}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}$. The $\rho$-metric of the latter satisfies
$\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}({\underline{x}},{\underline{y}})=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}(x_{i},y_{i})\;\leqslant\;\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,X_{i}(x_{i},y_{i})=(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})({\underline{x}},{\underline{y}}).$
Since $\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}$ is symmetric, we have $\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}\leqslant(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}$. The reverse inequality is more subtle: take a sequence of $n+1$ points ${\underline{z}}^{j}$, which varies from $\underline{x}$} to $\underline{y}$} by changing one coordinate at a time
${\underline{z}}^{j}=(y_{1},...,y_{j},x_{j+1},...,x_{n}),\qquad{\underline{z}}^{0}={\underline{x}},\;\,{\underline{z}}^{n}={\underline{y}}\qquad(j=0,...,n).$
and apply the triangular inequality
$(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}({\underline{x}},{\underline{y}})\;\leqslant\;\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}({\underline{z}}^{j-1},{\underline{z}}^{j}).$
Now, ${\underline{z}}^{j-1}$ and ${\underline{z}}^{j}$ only differ at the $j$-th coordinate ($x_{j}$ or $y_{j}$, respectively); restricting the domain of the ‘inf’ in the right term above to those sequences in $(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}$ where only the $j$-th coordinate changes, we get the $\delta$-metric $X_{j}^{\,{\scriptstyle{\wedge}}\,}$, and the inequality
$\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}({\underline{z}}^{j-1},{\underline{z}}^{j})\;\leqslant\;\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,X_{j}^{\,{\scriptstyle{\wedge}}\,}(x_{j},y_{j})=(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}_{j}\,X_{j})^{\,{\scriptstyle{\wedge}}\,}({\underline{x}},{\underline{y}}).$
(b) In fact
$\begin{array}[]{c}(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})^{\,{\scriptstyle{\vee}}\,}({\underline{x}},{\underline{y}})=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,X_{i}(x_{i},y_{i})){\,{\scriptstyle{\vee}}\,}(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,X_{i}(y_{i},x_{i})\\[5.0pt] \qquad\qquad\qquad\leqslant\;\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{i}\,(X_{i}(x_{i},y_{i}){\,{\scriptstyle{\vee}}\,}X_{i}(y_{i},x_{i}))=\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i}^{\,{\scriptstyle{\vee}}\,}({\underline{x}},{\underline{y}}).\end{array}$
Finally, point (c) is obvious. For (83), the inequality $v\langle a,b\rangle\leqslant v(a)+v(b)$ is obvious, and the other follows from the first remark in the proof of Theorem 5.5: given a partition for $a$ and one for $b$, by using a common refinement for both we get higher values. ∎
6.3. Corollary I (Reflective symmetrisation and products)
(a) For a finite family of $\rho$-metric spaces $X_{1},...,X_{n}$
(84) $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\wedge}}\,}\leqslant(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\wedge}}\,},\qquad(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\wedge}}\,}=\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i}^{\,{\scriptstyle{\wedge}}\,}\leqslant\;n(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\wedge}}\,}).$
(b) More particularly, for $\delta$-spaces
(85) $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\wedge}}\,}\leqslant(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\wedge}}\,}\leqslant\,(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\wedge}}\,}=\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i}^{\,{\scriptstyle{\wedge}}\,}\leqslant n(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\wedge}}\,}).$
All these symmetric $\delta$-metrics are Lipschitz-equivalent and induce the same topology: the functors $^{\,{\scriptstyle{\wedge}}\,}\colon\delta_{\infty}\mathsf{Mtr}\to\;!\delta_{\infty}\mathsf{Mtr}$ and $^{\,{\scriptstyle{\wedge}}\,}\colon\delta_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ preserve finite products and take tensor products to cartesian products.
Proof.
(a) To compare $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\wedge}}\,}$ and $(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\wedge}}\,}$, we note that
$(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\wedge}}\,})({\underline{x}},{\underline{y}})={\rm sup}_{i}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}(x_{i},y_{i})\leqslant{\rm sup}_{i}\,X_{i}(x_{i},y_{i})=(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})({\underline{x}},{\underline{y}}).$
Since $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}(X_{i}^{\,{\scriptstyle{\wedge}}\,})$ is symmetric, it follows that $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}(X_{i}^{\,{\scriptstyle{\wedge}}\,})\leqslant(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\wedge}}\,}$. The second part follows from (79).
(b) For $\delta$-spaces, the inequality $(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}\leqslant(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}$ follows from $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}\leqslant\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i}$, proved in (76). ∎
6.4. Corollary II (Coreflective symmetrisation and products)
(a) For a finite family of $\rho$-spaces $X_{1},...,X_{n}$
(86) $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\vee}}\,}=(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,},\qquad(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}\leqslant\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i}^{\,{\scriptstyle{\vee}}\,}\leqslant n(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\vee}}\,}).$
Therefore the functor $^{\,{\scriptstyle{\vee}}\,}\colon\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ preserves finite products (but takes arbitrary products to the box topology).
(b) More particularly, for $\delta$-spaces
(87) $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\vee}}\,}=(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}\leqslant(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}\leqslant\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}(X_{i}^{\,{\scriptstyle{\vee}}\,})\leqslant n(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\vee}}\,}).$
Again, these symmetric $\delta$-metrics are Lipschitz-equivalent and induce the same topology: the functor $^{\,{\scriptstyle{\vee}}\,}\colon\delta_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ preserves finite products, and takes tensor products to cartesian products.
(c) If the symmetrisations $X_{i}^{\,{\scriptstyle{\vee}}\,}$ of the $\rho$-spaces $X_{i}$ are positive
(88) $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\vee}}\,}=(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}\leqslant\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}(X_{i}^{\,{\scriptstyle{\vee}}\,})\leqslant n(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\vee}}\,}),$
and again these $\rho$-spaces are Lipschitz-equivalent (but $(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}$ need not be).
Proof.
(a) Coreflective symmetrisation is a right adjoint, and preserves products. The relation $(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}\leqslant\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i}^{\,{\scriptstyle{\vee}}\,}$ is proved in 6.2, and the last inequality in (86) is obvious, as in (77).
(b) As in Corollary I, the inequality $(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}\leqslant(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}$ follows from $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}\leqslant\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}X_{i}$, proved in (76) for $\delta$-spaces. We know that this does not extend to $\rho$-spaces, even in the symmetric case.
(c) If all $X_{i}^{\,{\scriptstyle{\vee}}\,}$ are positive, the inequality $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i}^{\,{\scriptstyle{\vee}}\,}\leqslant\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}(X_{i}^{\,{\scriptstyle{\vee}}\,})$ is a straightforward consequence of (76), and we already know that $(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}X_{i})^{\,{\scriptstyle{\vee}}\,}=(\raisebox{1.93747pt}{${\mbox{$\prod$}}$}(X_{i}^{\,{\scriptstyle{\vee}}\,})$. ∎
6.5. Ascent and descent
(a) Re-examining the elementary example of the Introduction, in 0.2, we consider the $\rho$-space $X=\delta\mathbb{R}\!\times\!\rho\mathbb{R}$, with the product $\rho$-metric $\rho((x,y),(x^{\prime},y^{\prime}))=\delta(x,x^{\prime}){\,{\scriptstyle{\vee}}\,}\rho(y,y^{\prime})$
(89) $\rho((x,y),(x^{\prime},y^{\prime}))\;=\;\begin{cases}(x^{\prime}-x){\,{\scriptstyle{\vee}}\,}(y^{\prime}-y),&\;\mbox{ if }\>x\leqslant x^{\prime},\\[3.0pt] \infty&otherwise.\end{cases}$
(b) A $\rho_{\infty}$-map $a=\langle a_{1},a_{2}\rangle\colon\delta\mathbb{I}\to X$
(90) ${a}$${x}$${y}\textstyle{{\scriptscriptstyle\bullet}}\textstyle{{\scriptscriptstyle\bullet}}$
has components $a_{1},a_{2}$ such that
- the path $a_{1}\colon\delta\mathbb{I}\to\delta\mathbb{R}$ is any increasing map $\mathbb{I}\to\mathbb{R}$ which is Lipschitz for the euclidean metric, as proved in 4.5(a),
- the path $a_{2}\colon\delta\mathbb{I}\to\rho\mathbb{R}$ satisfies the Lipschitz condition, examined in 4.6 and 5.1(d).
(c) Evaluation, total ascent, total descent and total variation have already been examined in 0.2, consistently with the definitions of 5.2.
6.6. Gravitational potential and metric
(a) The gravitational potential around the Earth, for a unit mass, can be expressed as
(91) $V(r)=-kr_{0}/r,\qquad\qquad\mbox{for }r\geqslant r_{0},$
where $r_{0}$ is radius of the Earth, $r\geqslant r_{0}$ is the distance from the centre of the Earth, and $k=GM/r_{0}$ is a positive constant; $G$ is the gravitational constant and $M$ is the mass of the Earth.
We rewrite it in the form
(92) $\Phi(x)=k-k/(1+x)=kx/(1+x),\qquad\;\;\mbox{for }x\geqslant 0,$
where $x=r/r_{0}-1\geqslant 0$ is the altitude, measured with respect to the Earth radius; we have added the constant $k$ so that the potential $\Phi$ annihilates at the Earth surface: $\Phi(0)=0$. The function $\Phi$ is increasing.
The associated (linear) $\rho$-metric $\rho(x,y)=\Phi(y)-\Phi(x)$ measures the gain of potential energy from altitude $x$ to altitude $y$, or the opposite of the work of the gravitational field in this transition.
(b) At small altitude, that is, for small $x$, one can use a linear approximation $\psi(x)=kx$ (which amounts to taking the scalar gravitational field constant), getting a $\rho$-metric
(93) $\rho^{\prime}(x,y)=\psi(y)-\psi(x)=k(y-x)\qquad\;\;\mbox{for }x,y\geqslant 0,$
which is linear and invariant up to (positive) translations.
(c) Consider now a spherical mass $M$ of radius $r_{0}$, centred at the origin of $\mathbb{R}^{3}$. Its gravitational potential, for a unit mass, can be expressed as
[…]
An interested reader can extend this to the gravitational (or electric) potential generated by a finite distribution of spherical masses (or charges), in the euclidean space.
6.7. Elastic potential and metric
(a) In the cartesian plane $xy$ a (linear) elastic field centred at the origin gives a potential function
(95) $\Phi(x,y)=\lambda(x^{2}+y^{2}),\qquad\mbox{for }(x,y)\in\mathbb{R}^{2},$
where $k=2\lambda>0$ is the elastic constant of the field: F(x,y)=k(x,y)F(x,y)=-k(x,y). The associated (linear) $\rho$-metric
(96) $\rho((x,y),(x^{\prime},y^{\prime}))=\Phi(x^{\prime},y^{\prime})-\Phi(x,y)=\lambda((x^{\prime 2}+y^{\prime 2})-(x^{2}+y^{2})),$
measures the gain of potential energy from position $(x,y)$ to $(x^{\prime},y^{\prime})$, that is the opposite of the work of the elastic field in this transition.
Letting $\rho_{E}\mathbb{R}$ have $\rho_{E}(x,x^{\prime})=x^{\prime 2}-x^{2}$, we are considering the linear $\rho$-space $(\lambda\rho_{E}\mathbb{R})^{\otimes 2}$.